### v5

# Auditing Human Decision-Making in High-Stakes Environments via Prescriptive AI: A Stress-Test on Real-Time Tactical Management

Abstract
High-stakes decision-making is often compromised by cognitive biases and outcome dependency. Current AI models typically mimic historical human behavior, inheriting these biases and limiting their utility for normative improvement. Here, we introduce a Prescriptive AI framework designed to audit, rather than automate, human judgment in real-time environments. By decoupling decision quality from stochastic outcomes, we quantify ”decision latency” and status quo bias in elite soccer management—a high-pressure adversarial domain. Analyzing 2018 FIFA World Cup data, our system exposes critical risk states, such as performance collapse following salient positive events (e.g., an assist), which human experts systematically overlook due to outcome bias. These findings demonstrate that interpretable auditing systems can reveal structural flaws in human reasoning that predictive models obscure. This approach establishes a paradigm for Human-AI interaction prioritizing epistemic accountability over predictive mimicry in safety-critical domains.
1 Introduction
Decision making under uncertainty in high-stakes environments constitutes a central challenge for human judgment. Settings such as clinical triage, operational risk management, and strategic resource allocation are characterized by irreversible actions, asymmetric consequences, and a critical dependence on timely and well-calibrated decisions. In such domains, decision errors are costly not only because of incorrect outcomes, but because they are often unjustified, delayed, or poorly aligned with underlying risk.
Despite these constraints, computational decision-support systems have been dominated by a fundamentally predictive paradigm. Models are typically optimized to replicate historical decision patterns, achieving statistical accuracy by mimicking prior human behavior. However, reliance on behavioral imitation is problematic: historical decisions often encode cognitive biases and institutional inertia. As a result, predictive systems create a structural mismatch between accuracy and decision quality, reaching a documented “predictive ceiling”—for instance, state-of-the-art substitution models plateau at approximately $70\%$ accuracy by merely cloning human choices (mohandas). Moreover, standard evaluation practices in stochastic environments suffer from outcome bias, judging decision quality by realized results rather than by the epistemic justification available at the time of choice (BaronHershey1988).
Importantly, this work does not aim to improve predictive accuracy, automate decision-making, or replace human expertise. Instead, it focuses on auditing the quality of human judgment under uncertainty by separating epistemic justification from stochastic outcomes. From this perspective, the limitations of predictive systems are not merely due to data insufficiency, but reflect a normative flaw inherent to supervised learning in agentic contexts. By mapping contextual inputs to historical decisions, such systems inevitably reproduce embedded cognitive biases, including status quo bias (reluctance to change strategy) and sunk cost fallacy (persistence in failing courses of action). Consequently, predictive models tend to validate conservative human behavior rather than reveal when deviation is epistemically warranted.
To address this limitation, we distinguish between Predictive AI, which forecasts events or replicates past decisions, and Prescriptive AI, defined here as a class of systems whose primary function is to audit, justify, and support human decision-making under uncertainty. This perspective aligns with decision-theoretic and epistemic accounts of agency, in which rational choice is defined as a coherent state transition at time $t$, independent of the stochastic outcome realized at time $t+n$ (vanBenthem2011). While prescriptive analytics and interpretable decision-support frameworks have been advocated in theory (Bertsimas2020; Lepenioti2020; Rudin2019), empirical demonstrations in realistic, adversarial environments remain scarce. Black-box models are particularly ill-suited for such auditing tasks due to automation bias and opacity (Parasuraman2010), motivating the need for systems that explicitly map observed states to epistemically justifiable actions.
In this work, we use elite-level soccer as a demanding natural laboratory to operationalize this prescriptive auditing paradigm. Substitution decisions share structural properties with many safety-critical contexts: they are time-sensitive, irreversible, and made under pervasive uncertainty. Existing approaches in this domain are limited either by insufficient temporal resolution—introducing exposure bias by favoring cumulative playing time over efficiency (Pappalardo2019; SchmidtLilloBustos2024)—or by a reliance on purely predictive modeling that fails to capture normative tactical necessity (wu; pamukkale).
To enable systematic auditing of human judgment, we propose a hybrid statistical–symbolic framework that overlays intrinsically interpretable fuzzy reasoning onto robust statistical signals. Rather than automating decisions, the system functions as an auditing layer, surfacing decision-relevant risk—such as performance decay or defensive liability—before it manifests as observable failure. Unlike descriptive fuzzy models (Marliere2017; zhou) or static ranking systems (salabun), the proposed approach operationalizes continuous, role-aware evaluation of decision states.
This work makes three contributions. First, we formalize a prescriptive auditing framework for evaluating human decision-making under uncertainty, in which symbolic reasoning contextualizes statistical evidence rather than replacing human agency. Second, we introduce a role-aware cumulative mean metric that eliminates play-time exposure bias, enabling the detection of intra-match performance deterioration. Third, we provide empirical validation in a high-stakes adversarial setting, demonstrating that the proposed auditing system reveals latent risk patterns systematically overlooked by both human experts and black-box predictive models.
2 Results
To evaluate the utility of prescriptive auditing in high-stakes environments, we contrast the proposed Fuzzy Control System (FCS) against state-of-the-art predictive baselines and subject it to a rigorous stress-test using real-world data from elite soccer.
2.1 Overcoming the Predictive Ceiling: Prescription vs. Prediction
Recent benchmarks in soccer analytics reveal a “predictive ceiling” for substitution modeling. mohandas analyzed 51,738 substitutions using supervised learning algorithms (Random Forest, SVM), achieving a maximum accuracy of $\approx 70\%$. This plateau suggests that nearly 30% of tactical decisions are driven by contextual nuances—such as fatigue thresholds or specific game plans—that remain opaque to black-box models.
Critically, supervised models in this domain suffer from a normative flaw: they are trained to mimic the human expert. If historical data encodes risk aversion or delayed reaction times, the ML model learns to replicate these biases. As detailed in Table 1, our proposed framework shifts the objective from minimizing prediction error to maximizing tactical utility, functioning as an auditor rather than an imitator.
Table 1: Structural Comparison: Supervised ML vs. Prescriptive Fuzzy System
Dimension Supervised ML Baseline (e.g., mohandas) Proposed Prescriptive System
Objective Minimize Prediction Error (Mimic Human) Maximize Tactical Utility (Audit Human)
Handling Uncertainty Probabilistic (Confidence Scores) Possibilistic (Membership Degrees)
Interpretability Low (Feature Importance/Black Box) High (Linguistic Rules)
Outcome Bias High (Dependent on realized results) Decoupled (State-based evaluation)
2.2 The Mathematical Mechanism: Eliminating Exposure Bias
A fundamental prerequisite for effective auditing is the ability to detect performance deterioration in real-time. Traditional metrics typically utilize cumulative sum formulations, where a player’s score monotonically increases with playing time regardless of minute-by-minute efficiency. Mathematically, these additive models act as low-pass filters, masking declining performance because a fatigued player performing poorly in the $80^{th}$ minute may still possess a higher total score than a high-impact substitute.
To resolve this Metric Exposure Bias, we reformulated the evaluation metric from a cumulative sum to a Cumulative Mean with Role-Aware Normalization ($P_{cum}$). By calculating the average contribution up to time $t$ and normalizing it against the historical distribution of the agent’s specific tactical role, our metric functions as a high-pass audit. This transformation allows the performance curve to reflect true temporal variability, exposing negative momentum and performance drops that are mathematically invisible in additive models. This methodological shift is what enables the system to detect the latent risks detailed in the following case studies.
2.3 Stress-Testing Decision Latency: The Brazil vs. Belgium Case
We applied the framework to the high-stakes elimination match between Brazil and Belgium (2018 FIFA World Cup) to observe decision latency and normative disagreement. The system outputs ($P_{final}$) were generated in 5-minute intervals.
Figure 1: Temporal evolution of Substitution Priority for Brazil. High values indicate critical inefficacy or risk. Note the divergence in the case of Fagner (Red Dashed Line) compared to executed substitutions.
2.3.1 Alignment and Early Detection
The system successfully captured decision-relevant deterioration prior to human intervention (Table 2).
  • Willian (45’): Priority $P_{final}=72.0$, ranked 4th. Rule R04 (Rapid Drop) flagged a sharp decline in technical actions in the final 20 minutes of the first half. Immediate substitution by the coach confirms offensive ineffectiveness, supported by media reports ESPN_2018.
  • Gabriel Jesus (58’): Priority $P_{final}=99.1$, 2nd highest. Rule R08 (Ineffective Forward) triggered by low cumulative score and no goals/assists. Post-match data: 12 touches in 58 minutes Epoca_2018, validating tactical isolation.
  • Paulinho (73’): Priority $P_{final}=93.1$. System detected ”Early Fatigue” and low participation. External analysis noted minimal offensive impact Gazeta_2018.
2.3.2 The “Fagner Paradox”: Counterfactual Auditing
The most significant finding is the disagreement regarding right-back Fagner. While the human decision-maker retained the player, the system assigned him Maximum Priority (100.0) from minute 45 to 85. This was not a prediction error, but a risk audit. The system aggregated consistently low technical performance with high defensive exposure against opponent Eden Hazard. The realized outcome—Fagner struggling in duels and receiving a yellow card at the 90th minute—validates the system’s “Critical Risk” diagnosis. A predictive ML model trained on this coach’s history would likely have predicted “No Substitution” (True Positive), thereby reinforcing the status quo bias that led to the vulnerability.
Table 2: Model Audit vs. Human Decision (Brazil vs. Belgium)
Player Analysis Slice($t$) $P_{\text{final}}$ Rank Diagnosis Real Decision
Willian 40’-45’ 72.0 4th Rapid Drop Substituted out
Gabriel Jesus 55’-60’ 99.1 2nd Ineffective Forward Substituted out
Fagner 45’-85’ 100.0 1st Critical Risk Not Substituted
Paulinho 65’-73’ 93.1 3rd Early Fatigue Substituted out
2.3.3 The “Lukaku Paradox”: Latency and Event Masking
Applying the system to the opponent (Belgium, Figure 2) highlights the impact of outcome bias on human decision-making.
Figure 2: Temporal evolution for Belgium. Lukaku reaches maximum priority (100.0) significantly earlier than his substitution.
Romelu Lukaku provided an assist at the 31st minute. While this single salient event seemingly justified his presence to the human manager, the system detected a severe drop in engagement immediately after.
  • Algorithmic Detection (35’): The system identified Lukaku as the #1 substitution candidate immediately post-assist, due to critically low volume (14 ball touches in 87 minutes).
  • Human Latency (+50 min): The coach waited until the 87th minute to substitute him.
This 50-minute latency demonstrates how prescriptive systems can decouple decision quality from “highlight moments,” ensuring consistent performance evaluation (Table 3).
Table 3: Comparative Analysis of Low-Volume Strikers
Player Context FCS Priority Human Response
G. Jesus (BRA) 12 Touches, 0 G/A High (#2) Substituted (58’)
R. Lukaku (BEL) 14 Touches, 1 Assist Max (#1) Delayed Sub (87’)
2.4 Systemic Validation: Longitudinal Patterns
Extending the analysis beyond single instances, we processed a representative subset of 25 World Cup matches—including the complete Brazilian campaign. This broader scope, comprising N = 9,152 decision windows, reveals structural regularities in how the system audits performance.
Technical Anchors vs. Chronic Inefficacy.
The system consistently distinguished between effective stability and passive stagnation. Key players like Coutinho and Marcelo functioned as “Technical Anchors,” maintaining priority scores near zero regardless of match time. Conversely, the system identified “Chronic Inefficacy” in players like Gabriel Jesus across multiple matches, where priority scores escalated without salient errors, supporting the hypothesis that the system detects latent risks invisible to event-based analysis.
Temporal Gap and Anticipation.
A systematic temporal gap was observed between the system’s “Critical” signal ($P_{final}>90$) and human intervention. In cases like Paulinho (vs. Serbia), the system flagged inefficacy several minutes prior to the substitution, quantifying the decision latency inherent in human processing under pressure.
Post-Entry Validation.
By tracking priority scores after a substitution, the system enables counterfactual assessment. Substitutes like Renato Augusto exhibited immediate monotonic decreases in priority (indicating high impact), whereas others (e.g., Fernandinho vs. Switzerland) maintained high risk scores, objectively characterizing the substitution as ineffective.
2.5 Boundary Conditions: The Limits of Performance Auditing
The system’s epistemic boundaries were rigorously tested by the case of Nacer Chadli (Belgium vs. Brazil), who was substituted at the 83rd minute due to an acute injury. This was the only substitution in the analyzed set not anticipated by the FCS ($P_{final}<50.0$).
Far from a systemic failure, this “false negative” provides crucial validation of the framework’s internal logic. Since the injury was a sudden, exogenous event unrelated to prior performance decay or observable fatigue, the system correctly refrained from flagging the player. This confirms that the high priority scores assigned in other cases (e.g., Lukaku, Jesus) were driven by genuine performance degradation, not by a generalized bias toward late-game substitutions. The framework effectively distinguishes between tactical necessity (predictable via data) and force majeure (epistemically inaccessible).
3 Discussion
This study leverages elite sports as a naturalistic laboratory to audit human judgment under uncertainty, revealing systematic divergences between normative risk assessment and expert human behavior. Our primary contribution to human decision science is the quantification of decision latency—the temporal gap between the statistical accumulation of risk and the human reaction to it. By deploying a system designed to decouple decision quality from stochastic outcomes, we provide empirical evidence that human experts in high-stakes environments are subject to structural cognitive biases—specifically status quo bias and outcome bias—which systematically delay necessary interventions.
To rigorize this auditing approach, we formalize the concept of Prescriptive AI:
Definition. Prescriptive AI is a class of intelligent systems whose primary function is to audit, justify, and support human decisions under uncertainty, rather than to automate actions or predict future outcomes.
Unlike standard predictive analytics, this paradigm is defined by five normative objectives:
  • A primary function of decision auditing rather than decision automation;
  • Explicit reasoning under uncertainty and asymmetric risk;
  • Intrinsic interpretability as a structural requirement, not a post-hoc feature (Rudin2019);
  • Contestable recommendations that preserve human agency;
  • Evaluation criteria that decouple state assessment from outcome realization, explicitly avoiding outcome bias (BaronHershey1988).
These properties impose normative constraints that exclude many systems commonly labeled as prescriptive. While the framework itself is model-agnostic, the system presented here serves as a concrete instantiation of these principles. By applying this approach, we show that the limitation of human experts is not a failure of strategic intent, but a failure of real-time information processing: humans struggle to detect non-salient performance decay (the “boiling frog” effect) until a catastrophic or highly salient event occurs.
3.1 Behavioral Audit: Quantifying Cognitive Biases
The empirical disparities between the prescriptive system and human agents are not random errors, but operationalizations of well-documented cognitive heuristics:
The “Fagner Paradox” as Status Quo Bias.
The disagreement regarding the defensive player Fagner illustrates the status quo bias (or omission bias) under pressure. While the system identified a “Critical Risk” state ($P_{final}=100.0$) based on sustained defensive exposure, the human manager chose inaction. This behavior reflects an asymmetric preference for maintaining the current state, requiring substantially stronger evidence to justify change than to justify persistence. Given the informational state available at time $t$, this inertia is normatively difficult to justify.
The “Lukaku Paradox” as Salience Masking and Outcome Bias.
The case of the opposing striker demonstrates how outcome bias distorts real-time evaluation. A single salient positive event (an assist) masked a severe and prolonged deterioration in engagement (over $50$ minutes of inactivity). We term this phenomenon Salience Masking: the overweighting of vivid, recent outcomes at the expense of cumulative base-rate information. The prescriptive system—immune to the “halo effect” of the assist—identified the decision to maintain the player as a high-risk error well before the human agent reacted.
3.2 Implications for Human Decision-Making
These findings extend classical accounts of Bounded Rationality in dynamic environments. We argue that decision latency is not merely a domain-specific artifact, but a domain-general construct reflecting the cognitive cost required to override heuristic inertia in the presence of uncertainty and time pressure.
  • Filtering and Blind Spots: Real-time tactical management requires processing high-dimensional, noisy data streams. To preserve strategic coherence, human agents apply low-pass cognitive filters that attenuate short-term fluctuations. Our results indicate that this adaptive filtering introduces a structural blind spot for gradual but systematic performance decay.
  • AI as Cognitive Orthotics: Accordingly, the role of Prescriptive AI shifts from an “oracle” that predicts future states to a cognitive orthotic—a support structure designed to compensate for specific, predictable human limitations. By surfacing latent risk through interpretable signals (e.g., Rule R04: Rapid Drop), the system confronts decision-makers with evidence their heuristic reasoning would otherwise suppress.
3.3 Normative Foundations: Decision-Making as Epistemic State Transition
The validity of this auditing approach rests on a dynamic conception of rationality, grounded in the logical dynamics framework of vanBenthem2011. Under this view, rationality is a property of the epistemic update at the moment of commitment ($t$), rather than of the stochastic realization of outcomes at a later time ($t+n$). If decision quality is defined prior to outcome realization, then systems designed to support human judgment must operate strictly on the informational state available at the moment of action. This perspective justifies our rejection of outcome-based evaluation metrics: a decision is not “correct” because the team won, but because it constituted a normatively coherent response to the risk state at time $t$.
Naturally, any prescriptive audit is contingent on the chosen normative model of risk. Alternative normative assumptions would yield different—but still auditable—recommendations, without undermining the core principle of decoupling decision quality from outcome realization.
3.4 Ethical Accountability and Epistemic Trust
Finally, we argue that in high-stakes human–AI teaming, interpretability is a pre-condition for legitimacy. Auditing human judgment requires more than predictive accuracy; it requires explanations that enable decision-makers to validate machine assessments against their own reasoning.
Resolving Epistemic Opacity.
In domains subject to intense public scrutiny—such as elite sports, clinical triage, or emergency operations—opaque “black-box” models pose a substantial epistemic risk. When an algorithm recommends a counter-intuitive action (e.g., substituting a star player), a lack of transparency forces stakeholders to choose between blind faith and outright rejection. By providing a transparent audit trail in which each recommendation is traceable to explicit linguistic rules, the proposed fuzzy framework reduces epistemic asymmetry and transforms the interaction from blind obedience into informed deliberation.
The Audit Trail as Institutional Defense.
Beyond decision support, prescriptive systems serve a protective institutional function. In environments vulnerable to suspicion—whether due to financial incentives, reputational risk, or public accountability—a documented audit trail allows decision-makers to demonstrate that their actions (or inaction) were consistent with a normative risk assessment given the information available at time $t$. In this sense, the system functions not merely as a tactical aid, but as an objective witness to the rationality of the decision process.
Ultimately, this work advances prescriptiveness as an ethical design principle. In contexts where decisions are public, consequential, and irreversible, only transparent, auditable, and human-centered systems can achieve the level of trust required to augment—rather than replace—human agency.
4 Methodology
4.1 Baseline Limitations and Methodological Motivation
The academic literature on soccer analytics provides the theoretical foundation for this work, situated at the intersection of performance evaluation and tactical decision modeling.
A central methodological challenge in this domain is quantifying player performance. The PlayeRank framework, proposed by Pappalardo2019, defines a multidimensional, role-aware metric that combines individual technical actions with influence on the passing network, validated against professional scout assessments. The authors explicitly aim to reduce biases related to playing time, normalizing performance vectors by the number of actions to “avoid biases due to play time.”
However, although normalization occurs at the level of the action vector, the final match score is computed as a weighted cumulative sum of contributions. When inspected over the temporal progression of a match, this formulation tends to produce strictly increasing trajectories, effectively reintroducing a play-time exposure bias that favors players who remain longer on the field, even if their per-minute contribution rates are similar. This monotonic behavior introduces a methodological limitation for temporal or comparative analyses. Each additional action contributes positively to the cumulative sum—regardless of qualitative impact—so players with identical per-minute performance profiles diverge in total score purely as a function of playing time. While this approach is suitable for match-level summaries, it becomes problematic when the goal is to characterize relative influence or detect shifts in performance dynamics.
The strictly increasing nature of the cumulative score also prevents the representation of performance decay, momentum reversals, or tactical suppression: a player whose influence declines due to fatigue or tactical adjustments will still exhibit an artificially monotonic performance trajectory. Therefore, while PlayeRank’s cumulative formulation is adequate for full-match evaluation, it is insufficient for temporal segmentation. A temporally meaningful metric must reflect evolving game states, substitutions, and local contributions, rather than indiscriminately accumulating actions.
Recent work supports this perspective. SchmidtLilloBustos2024 demonstrate that player influence fluctuates meaningfully over a match due to tactical reconfiguration, opponent pressure, and physical factors, advocating for temporal windowing to capture such dynamics. Following this insight, the present work also adopts 5-minute temporal slices, but advances the methodology by adding a role-aware percentile cumulative mean across slices. This modification preserves interpretability while enabling direct comparability between players with different playing times, eliminating residual exposure bias and allowing the performance curve to reflect true temporal variability.
A second methodological challenge concerns the interpretation of imprecise or uncertain information in decision-making contexts. The literature highlights fuzzy logic as an effective tool for this purpose. For instance, Huarachi-Macuri et al. LACCEI2023 demonstrate that fuzzy systems are suitable for transforming quantitative descriptors into interpretable linguistic assessments (e.g., Stamina, Agility). Within tactical modeling, Marliere2017 uses fuzzy control systems to arbitrate between tactical states under uncertainty. These contributions motivate the role of fuzzy logic in the present system: integrating performance indicators and contextual factors into a Substitution Priority score in a manner compatible with the inherent uncertainty of in-game decision making.
The decision to make substitutions in soccer remains an intrinsically human domain, hardly replicable by machine learning (ML) algorithms. While ML can analyze historical performance data, a coach’s choice involves a complex assessment of contextual, tactical, and psychological factors that are not easily quantifiable—such as team morale, perceived fatigue, reading the opponent’s strategy, and game momentum. Attempts to predict substitutions using ML, while informative, reveal a clear ceiling in their ability to capture this complexity. A recent study by mohandas, for example, analyzed a large dataset of 51,738 substitutions using multiple algorithms, including Random Forest, SVM, and Decision Trees. Even with this wealth of data, the best-performing model (Random Forest) achieved a maximum accuracy of just over 70%. This gap of approximately 30% is significant, suggesting that the final decision is not purely predictive but rather adaptive, relying on subjective human judgment that historical data alone cannot capture. Consequently, models attempting to predict the optimal substitution are conceptually limited. Alternative approaches, such as fuzzy logic-based decision support systems, appear more suitable, as they are designed not to replace but to support the coach’s judgment, managing the inherent uncertainty and qualitative factors of the match.
Finally, the choice of an intrinsically interpretable methodology is not merely a design preference, but a methodological necessity in high-stakes decision-making contexts. As explicitly argued by Rudin2019, post-hoc explanations of black-box models are fundamentally inadequate when decisions have significant consequences, as they obscure the true decision logic and may provide misleading justifications. In domains such as elite soccer, where substitution decisions carry substantial sporting and financial impact, decision support systems must expose their reasoning process in a transparent and auditable manner. This perspective reinforces the limitations of purely predictive machine learning approaches discussed above and provides a principled justification for abandoning black-box optimization in favor of symbolic reasoning. Accordingly, the present work adopts fuzzy logic as the core inference mechanism—rather than as an explanatory add-on—ensuring that substitution recommendations are interpretable, accountable, and directly aligned with domain knowledge and human tactical judgment (Rudin2019).
This perspective finds foundational support in the framework of logical dynamics proposed by vanBenthem2011, which conceptualizes rational agency as a continuous process of information update rather than static optimization. In this context, player substitutions serve as a particularly clear instance of epistemic actions (vanBenthem2007): they do not merely alter the physical parameters of the match, but fundamentally restructure the space of plausible future game states under severe informational constraints. At the moment of decision, the coach operates at the boundary between “hard facts” (e.g., scoreline, remaining time, substitution limits) and “soft information” (e.g., perceived fatigue, tactical momentum). Once executed, a substitution irreversibly fixes certain constraints while reweighting the plausibilities regarding future play. Consequently, decision quality cannot be assessed solely by ex-post outcomes, but by whether the action constituted a rational epistemic update given the information available at time $t$. This framing aligns the proposed system with a dynamic view of rationality—where correctness is understood as the capacity for informed correction under uncertainty—establishing its role not as a predictor of future success, but as an auditor of the epistemic transition induced by the substitution.
Crucially, the proposed approach is hybrid in a functional rather than predictive sense. The statistical components of the system are not employed to forecast future match outcomes or to optimize predictive accuracy. Instead, they serve exclusively to construct a temporally normalized, role-aware representation of the current decision state. The symbolic layer does not refine or correct predictions; it operates as an external normative auditor, evaluating whether continued action is justified given the evidence available at time $t$. In this architecture, hybridization exists to support interpretability, decision auditing, and accountability—not predictive performance.
4.2 Description of the Problem
4.2.1 Problem Formulation
Tactical decision-making in elite soccer operates under high uncertainty and significant financial stakes GE_PremierLeague; ESPN_Mundial. Despite the proliferation of granular data, the specific process of substitution decisions remains predominantly intuitive or reliant on descriptive statistics that fail to capture real-time performance decay. This creates two distinct problems: the Metric Exposure Bias and the Predictive Ceiling.
The Metric Exposure Bias: Existing frameworks for player evaluation, such as the original PlayeRank Pappalardo2019, typically utilize cumulative sum formulations. While effective for post-match rankings, these metrics introduce a temporal bias during live games: a player’s score monotonically increases with playing time, regardless of their minute-by-minute efficiency. This mathematical structure masks declining performance, as a fatigued player performing poorly in the $80^{th}$ minute may still have a higher total score than a high-impact substitute, rendering standard metrics insufficient for real-time substitution decisions.
The Predictive Ceiling of Machine Learning: Current computational approaches to substitutions rely heavily on Supervised Machine Learning (SML) models trained on historical data. As noted by mohandas, these models achieve a prediction accuracy plateau of approximately 70%. This ”ceiling” exists because SML models are designed to mimic human behavior, thereby learning and replicating the cognitive biases of coaches (e.g., status quo bias or delaying defensive changes). Consequently, these models validate historical decisions rather than optimizing future outcomes, failing to identify necessary substitutions that deviate from conservative human norms.
Therefore, the core problem this study addresses is the lack of an objective, prescriptive auditing tool that can quantify intra-match performance decay without exposure bias and signal tactical risks independently of historical human tendencies.
4.2.2 Research Hypotheses
To address the formulated problem, this study tests the following hypotheses:
  • H1 (Metric Hypothesis): Reformulating performance evaluation from a cumulative sum to a Cumulative Mean with Role-Aware Normalization effectively eliminates exposure bias, allowing the system to detect negative momentum and performance drops that are mathematically invisible in additive models.
  • H2 (Prescriptive Auditing Hypothesis): A prescriptive decision support system, acting as an objective auditor of performance and contextual variables (e.g., performance trends, fatigue, disciplinary exposure), can identify high-risk tactical scenarios—such as latent defensive liabilities—earlier and more reliably than human intuition or data-driven predictive models trained on historical behavior.
  • H3 (Contextual Modulation Hypothesis): The modulation of disciplinary risk by tactical role (e.g., weighting a yellow card more heavily for a defender than a forward) significantly alters the substitution priority score, aligning the system’s output with expert tactical consensus in critical defensive scenarios.
  • H4 (Prescriptive AI Formalization Hypothesis): The empirical validation of a prescriptive, auditing-based decision support system in a high-stakes environment supports the formalization of a distinct Prescriptive AI paradigm, in which the system objective is not to predict human decisions, but to support normative, risk-aware judgment through explicit reasoning and accountability.
4.3 Dataset
The database selected for this project is the ”Soccer match event dataset”, a detailed public repository of soccer match events (Pappalardo2020). The choice of this dataset is based on its high granularity. The dataset details player-level actions and aggregated performance metrics, allowing for the modeling of in-game situations, which is an essential requirement for our system. We used the complete dataset available on Kaggle (Pappalardo2020), which comprises 27 interrelated CSV tables. They cover seven main competitions and can be grouped into five logical categories:
  • Match Data (matches_*.csv): Contains information about the games, such as dates, lineups, substitutions, results, and tactical formations.
  • Event Data (events_*.csv): The core of the dataset, recording millions of individual actions on the field.
  • Entity Data (players.csv, teams.csv, etc.): Dimensional tables with demographic and static data.
  • Performance Metrics (playerank.csv): A pre-processed file that provides the playerankScore, a multidimensional and role-aware performance evaluation metric.
  • Dictionaries (tags2name.csv, eventid2name.csv): Metadata that translate event and tag IDs into readable descriptions (e.g., Tag 1702 = ’yellow_card’).
The choice of playerankScore as our main ”Performance” input is a central methodological decision. As proposed by Pappalardo2019, the PlayeRank framework was developed to solve the absence of a consolidated and universally accepted metric for evaluating player performance. The playerankScore is a metric derived from millions of game events that, according to the authors, surpasses other metrics when compared with assessments from professional scouts. Therefore, instead of trying to model performance from raw events, we adopted the playerankScore as an already validated and academically robust representation of a player’s performance in a match.
4.4 Data Integration and Pre-processing
The pre-processing pipeline transforms heterogeneous soccer event logs into a unified temporal representation of player performance suitable for fuzzy inference. The process unfolds in three sequential stages: temporal performance computation, contextual and demographic enrichment, and role-aware normalization. Each stage progressively increases the dataset’s semantic density and interpretability.
The objective of this pipeline is to transform the raw multi-source event data (events, matches, players) into a single structured and temporal dataset tailored for the Fuzzy Logic Decision System. Unlike the original PlayeRank framework Pappalardo2019, which aggregates player performance per match, our system requires a fine-grained temporal perspective of the player’s performance evolution within a match. To achieve this, we implemented a three-stage processing pipeline that generates a dense temporal dataset, where each record represents the state of a player in a 5-minute (300-second) interval of play.
4.4.1 Stage 1: Temporal Performance Metric Generation
The first stage (playerankdatasetV4.py) processes the raw event logs (events_*.csv) and converts each action’s timestamp (eventSec, matchPeriod) into an absolute measure of seconds from the start of the match (total_seconds). The match timeline is then discretized into 5-minute intervals, associating each event with its corresponding temporal slice. The following variables are created in this stage:
  • matchId, teamId, playerId: extracted directly from the event logs, providing the hierarchical identifiers for each observation.
  • cartao_amarelo: a binary state indicator (using expanding().max()) equal to 1 if the player received a yellow card at any point up to the current slice.
Within each interval, two complementary components are calculated:
  • Technical Score (score_tecnico_fatia): the aggregation of event-based actions (passes, shots, duels, interceptions) using the same weights of the original PlayeRank framework. Events are mapped to technical dimensions and normalized within the team context.
  • Network Score (score_rede_fatia): computed using a directed pass graph per team and interval. Following the original PlayeRank framework, the Eigenvector Centrality of each player is used to quantify their structural influence on ball circulation.
These two components are linearly combined to form the primary performance indicator for each temporal slice:
$\mathrm{playerank\_fatia\_raw}=(1-\alpha_{\mathrm{net}})\cdot\mathrm{score\_tecnico\_fatia}+\alpha_{\mathrm{net}}\cdot\mathrm{score\_rede\_fatia}$
where $\alpha_{\mathrm{net}}=0.2$ controls the relative contribution of the network-based component. This configuration was selected to balance structural influence and technical efficiency, aligning with the sensitivity ranges reported in Pappalardo2019.
Subsequently, to ensure metric interpretability independent of absolute match intensity, the raw score is normalized against the historical distribution of the agent’s specific tactical role. This transformation converts the unbounded scalar playerank_fatia_raw into the Role-Aware Percentile (playerank_fatia_percentil), ensuring that a ”High” score represents the same statistical rarity for a Defender as it does for a Forward.
Cumulative Metric Redefinition.
To capture performance evolution through time, we define the cumulative mean score:
$\mathrm{playerank\_acumulativo\_media\_percentil}\ \ \textbf{=}\ \ \frac{1}{t}\sum_{i=1}^{t}\mathrm{playerank\_fatia\_percentil}^{(i)}$
This cumulative formulation departs from the additive model of the original PlayeRank, ensuring that longer playing times do not artificially inflate the total performance. The metric represents the player’s average contribution up to time $t$, maintaining comparability among players with different play durations.
4.4.2 Stage 2: Contextual and Demographic Enrichment
The second stage (expandirdataset.py) enhances the temporal dataset with contextual, situational, and demographic variables by joining multiple data sources.
  • player_age: calculated by merging player metadata (players.csv) with match dates and computing the player’s age on match day.
  • player_position: extracted from the player metadata, representing the tactical role (goalkeeper, defender, midfielder, forward).
  • goals_scored: identified from events where eventName = "Shot" and subEventName = "Goal" and accumulated over time.
  • assists: extracted using Wyscout tag id = 302 and accumulated over time using groupby([’matchId’, ’playerId’]).cumsum().
  • momentum_rate: a measure of short-term performance trend, calculated as the difference between the current($t$) and the previous(t1t-1) 5-minute playerank_acumulativo_media_percentil slices.
This enrichment process yields a semantically rich dataset combining technical, contextual, and demographic aspects for each player’s temporal trace.
4.4.3 Stage 3: Final Cleaning and Role-Aware Normalization
The final stage (limpezafinal.py) standardizes the temporal data and recalculates cumulative metrics for stability and accuracy. The following operations are performed:
  • removal of out-of-field periods: Periods where it was identified that the player was substituted, or entered as a substitute, were removed to ensure that the metric remained accurate, not applied to goalkeepers.
  • Temporal Harmonization: The temporal variables are standardized. Tempo_Partida is set to represent the *end* of each 5-minute slice (e.g., 5, 10, 15…), and minutes_played is recalculated using cumcount() to ensure a precise cumulative sum of on-field time.
  • Recalculation of Cumulative Means: The variable playerank_acumulativo_media_raw is recalculated using a robust expanding().mean() to ensure that averages reflect the final, harmonized temporal trace.
  • Percentile Normalization: two percentile-based variables are computed within each (matchId, position) group. Note: this normalization uses the raw position column (from the original data source) to group players by tactical role, ensuring methodological consistency:
    $\mathrm{playerank\_fatia\_percentil}$ $=\mathrm{rank}_{\mathrm{pct}}(\mathrm{playerank\_fatia\_raw})$
    $\mathrm{playerank\_acumulativo\_media\_percentil}$ $=\mathrm{rank}_{\mathrm{pct}}(\mathrm{playerank\_acumulativo\_media\_raw})$
This role-aware normalization ensures that performance is assessed relative to tactical peers, preserving fairness and interpretability in subsequent fuzzy inference.
Final Dataset Variables.
The resulting dataset (Dataset_limpo_final.csv) contains the following variables:
  • Identifiers: matchId, playerId, teamId.
  • Temporal Dimensions: Tempo_Partida, minutes_played.
  • Performance Metrics: playerank_fatia_raw, playerank_acumulativo_media_raw, score_tecnico_fatia, score_rede_fatia.
  • Normalized Metrics: playerank_fatia_percentil,
    playerank_acumulativo_media_percentil.
  • Contextual Factors: momentum_rate, cartao_amarelo.
  • Demographics: player_age, player_position.
  • Offensive Contributions: goals_scored, assists.
  • Auxiliary: position, a placeholder variable maintained for compatibility.
Each record corresponds to a 5-minute temporal snapshot of an individual player’s on-field performance, integrating technical, contextual, and demographic information in a consistent temporal framework.
4.4.4 Data Validation
Before integration into the fuzzy system, consistency checks were performed. The recalculation of playerank_acumulativo_media_raw in the final stage (Stage 3) ensures that all cumulative metrics are based on the robust expanding().mean() operator applied to the complete and harmonized temporal trace of each player. This guarantees that all cumulative scores represent valid in-game performance dynamics and that the fuzzy system operates on stable, correctly aggregated data. The experimental design yields high statistical density: the inference engine executes approximately 360 distinct decision audits (2 teams × 10 outfield players × 18 temporal windows). This granular approach allows us to evaluate the system’s stability across diverse game states (winning/losing, high/low fatigue) within a controlled adversarial environment, ensuring internal validity beyond a simple match-outcome correlation
4.5 Exploratory Data Analysis
The Exploratory Data Analysis (EDA) was conducted on the final integrated dataset, consisting of 805,146 temporal observations, 3,035 unique players, and 1,941 matches. The analysis aimed to (i) characterize the statistical properties of the input variables used in the fuzzy inference system, (ii) validate soccer-specific behavioral hypotheses (e.g., positional risk asymmetry), and (iii) provide empirical grounding for the design of the fuzzy universes of discourse.
4.5.1 Descriptive Overview
The dataset preserves a balanced representation of real-world soccer demographics and tactical structures. Table 4 summarizes the player distribution and key descriptive statistics.
Table 4: Dataset Summary and Descriptive Statistics
Category Value Metric Notes
Total observations 805,146 intervals 5-minute resolution
Unique players 3,035 across 7 competitions
Matches analyzed 1,941 complete event coverage
Mean/Median performance 0.501 std = 0.289 centered around 0.5
(playerank...percentil) 0.501 symmetrical distribution
Player positions 4 roles role-aware normalization applied
   Goalkeeper 233 (5.8%)
   Defender 1,032 (25.8%)
   Midfielder 1,120 (28.0%)
   Forward 650 (16.3%)
This composition closely mirrors elite-level soccer distributions, where midfielders and defenders together comprise over half of the active roster, reflecting the denser tactical occupation of central zones.
4.5.2 Univariate Distributions
Player Age.
The age distribution (Figure 3) follows an approximately median centered around 26.0 years, with a range between 15.0 and 45.0 years. The kurtosis (2.91) and near-zero skewness confirm symmetry. This validates the fuzzy partitioning into Young, Peak, and Veteran sets, with smooth transitions around the mean.
Figure 3: Distribution of player ages in the dataset.
Performance (playerank_acumulativo_media_percentil)
The per-slice performance scores exhibit a bounded, quasi-normal distribution centered around 0.5, with a standard deviation of 0.289. These characteristics make the metric well-suited for fuzzy modeling, where zero-centered or bounded scales allow intuitive linguistic partitioning (Low, Medium, High) without rescaling distortions.
4.5.3 Role-Based Event Distributions
To further interpret positional behavior, the dataset was aggregated by player position to estimate empirical probabilities of key in-game events (goals, assists, and yellow cards).
The empirical distributions align with established tactical expectations:
  • Goal Rate: Forwards lead with a 19.8% probability of scoring per match, compared to 7.7% for midfielders and 3.4% for defenders. Goalkeepers are near zero, validating the offensive gradient embedded in the player_position variable.
  • Assist Rate: Forwards exhibit the highest assist probability (7.5%), reflecting their creative and distributive role in ball progression. Midfielders follow closely (6.8%), while defenders contribute the least(3.6%).
  • Disciplinary Risk: Defenders show the highest yellow card rate (17.2%), followed by midfielders (14.8%) and forwards (10.1%), reinforcing the need for contextual fuzzy rules that modulate card risk according to tactical role.
4.5.4 Variable Definition, Independence and Correlations
Variable Definition and Selection
From the final integrated dataset (Dataset_limpo _final.csv), nine variables were selected for exploratory analysis and subsequent use as inputs in the fuzzy decision model. These variables represent the main dimensions of in-game player state: performance, fatigue, disciplinary risk, offensive contribution, and tactical context. Table 5 summarizes the selected features, their measurement scale, and their conceptual role within the system.
Table 5: Variables selected for exploratory analysis and fuzzy modeling.
Dimension Variable (ID) Mathematical Set Description and Relevance Technical Performance P_cum $0,1]\subset\mathbb{R}$ Cumulative performance percentile per match and position; main indicator of global efficiency. Fatigue Min_played $0,100]\subset\mathbb{R}$ Total minutes played at each interval; proxy for physical wear. Disciplinary Risk Card_Y $\{0,1\}\subset\mathbb{Z}$ Indicates if player received a yellow card; higher risk increases substitution priority. Form Trend Momentum $1,1]\subset\mathbb{R}$ Short-term rate of change in performance (3-interval moving average). Negative values denote declining form. Contextual Fatigue (Age) Age $15,45]\subset\mathbb{N}$ Player’s age on match day; modulates fatigue effects. Tactical Role Position $Categorical$ Role label; used for role-aware normalization. Offensive Contribution (Goals) Goals $\mathbb{N}$ Cumulative goals scored; high values indicate offensive importance. Offensive Contribution (Assists) Assists $\mathbb{N}$ Cumulative assists; complements goal contribution.
Labels: P_cum = playerank_acumulativo_media_percentil, Min_played = minutes_played, Card_Y = cartao_amarelo, Momentum = momentum_rate, Age = player_age, Position = player_position, Goals = goals_scored, Assists = assists,
These variables were derived directly from the processed PlayeRank-based dataset, ensuring temporal coherence across 5-minute intervals. The combination of continuous, discrete, and categorical features provides a multidimensional view of player state suitable for fuzzy inference. In the next subsection, we assess the independence and correlation among these variables to confirm their suitability for use as fuzzy inputs.
Spearman Correlation Matrix
Figure 4 reports the Spearman correlation matrix computed over all input variables adopted in the fuzzy inference system. The pairwise correlations are consistently weak, with magnitudes predominantly below $|\rho|<0.20$. Such a pattern indicates that the variables do not exhibit meaningful monotonic dependence.
Figure 4: Spearman correlation matrix among the fuzzy-system input variables. The weak correlations indicate low redundancy and support the assumption of variable independence in fuzzy inference.
This low degree of association is desirable in the context of fuzzy inference. Since the rule base relies on linguistic terms and membership functions that encode distinct semantic dimensions of player performance, strongly correlated variables would introduce redundancy and reduce the discriminative power of the fuzzy rules. Conversely, the weak correlations observed here suggest that each variable contributes independent information to the inference process, thereby supporting a more expressive and interpretable fuzzy model.
4.6 Fuzzy Control System
Based on the theoretical foundation and the need for dynamic real-time evaluation, we designed the Fuzzy Control System (FCS) to function as a Contextual Modifier. Unlike traditional systems that calculate a raw output from zero, this architecture calculates a correction factor applied to a baseline performance metric. The system was implemented using the scikit-fuzzy library in Python, utilizing the Mamdani architecture.
4.6.1 Definition of Fuzzy Variables
The system architecture has been expanded to process high-dimensional match data. It comprises eight input variables (antecedents) and one output variable (consequent). The universes of discourse were calibrated based on the range of values observed in the dataset and domain constraints (e.g., match duration).
Antecedents (Inputs)
  1. 1.
    Cumulative Performance (P_cum): Represents the player’s average percentile performance up to the current moment.
    • Universe: [0.0, 1.0].
    • Sets: VeryLow, Low, Medium, High, VeryHigh.
  2. 2.
    Momentum (Momentum): Rate of change in performance, indicating if the player is improving or declining.
    • Universe: [-1.0, 1.0].
    • Sets: Falling, Stable, Rising.
  3. 3.
    Fatigue (Min_played): Minutes played in the match.
    • Universe: [0, 100] minutes.
    • Sets: Low (0-45’), Medium (40-80’), High (70-100’).
  4. 4.
    Age (Age): Player’s chronological age.
    • Universe: [15, 45] years.
    • Sets: Young, Peak, Veteran.
  5. 5.
    Match Events (Categorical/Integer):
    • Card_Y: Yellow card status [0, 1] (Yes).
    • Goals: Goals scored [0, 10] (None, Some, Many).
    • Assists: Assists provided [0, 10] (None, Some, Many).
  6. 6.
    Positional Context (Binary):
    • Variables: is_Defender, is_Midfielder, is_Forward.
    • Sets: Yes (indicating the player’s role to activate specific rules).
Consequent (Output)
  • Modifier Value (Modifier_Value): A correction factor ranging from negative (protection/priority reduction) to positive (urgency/priority increase).
    • Universe: [-100, 100].
    • Granularity: 9 sets ranging from VLN (Very Large Negative, -100) to LP_70 (Large Positive, +70).
    • Zero represents no adjustment.
4.6.2 Membership Functions
The membership functions (MFs) mix Trapezoidal (trapmf) and Triangular (trimf) shapes to capture specific tactical thresholds. Key definitions from the implementation include:
  • P_cum: Low uses a trapezoid [0, 0, 0.10, 0.35] to capture distinctively poor performance, while High starts at 0.65.
  • Momentum: Falling is defined strictly in the negative range [-1.0, -1.0, -0.03, -0.01], ensuring only genuine performance drops trigger the logic.
  • Fatigue: Overlapping sets allow smooth transitions. High fatigue begins notably at 70 minutes, aligning with common substitution windows.
  • Contextual Events: Variables like is_Defender or Card_Y use pseudo-binary trapezoids (e.g., [0.5, 1, 1, 1.5]) to function as logical switches within the fuzzy inference engine.
4.6.3 Rule Base
The rule base was streamlined to remove noise and focus on Intensifying or attenuating the player’s performance using other contextual events as variables. The system avoids oscillation by ignoring minor fluctuations and focuses on critical states. The logic integrates position and stats directly.
Table 6: Key Fuzzy Rules for Substitution Priority
ID Rule Logic R01 – Untouchable Star IF (P_cum is High/VeryHigh) THEN (Modifier = Negative). R02a – Critical Fatigue IF (P_cum Low/VeryLow) AND (Min High) THEN (Modifier = Large Positive). R02b – Early Fatigue IF (P_cum Low/VeryLow) AND (Min Medium) THEN (Modifier = Medium Positive). R03 – Defensive Risk IF (Defender AND YellowCard Yes) THEN (Modifier = Medium Positive). R04 – Rapid Decline IF (P_cum Low/VeryLow AND Momentum Falling) THEN (Modifier = Large Positive). R07 – Positive Momentum IF (P_cum High/VeryHigh AND Momentum Rising) THEN (Modifier = Medium Negative). R08 – Ineffective Forward IF (Forward AND P_cum Low AND Goals None) THEN (Modifier = Large Positive). R09 – Striker Under Pressure (1) IF (Forward AND P_cum Low/Med AND Momentum Falling) THEN (Modifier = VeryLarge Positive). (2) IF (Forward AND P_cum Low/Med AND Min High) THEN (Modifier = VeryLarge Positive). R10 – Invisible Playmaker IF (Midfielder AND P_cum Low AND Assists None) THEN (Modifier = Med–Large Positive). R11 – Creator Bonus IF (Assists Some OR Goals Some) THEN (Modifier = Medium Negative). IF (Assists Many OR Goals Many) THEN (Modifier = Large Negative). R12 – Veteran Fatigue IF (Age Veteran AND Min High) THEN (Modifier = Medium Positive). R13 – Young Talent Protection IF (Age Young AND (Goals Some/Many OR Assists Some/Many)) THEN (Modifier = VeryLarge Negative). R14 – Goal Protection IF (Goals Some) THEN (Modifier = Medium Negative). IF (Goals Many) THEN (Modifier = Large Negative). R15 – Neutral State IF (P_cum Medium AND Momentum Stable) THEN (Modifier = Zero).
4.6.4 Inference and Final Calculation
The system employs a hybrid calculation model. The Fuzzy Inference System (FIS) computes the Modifier_Value using the Centroid method. However, this value is not the final priority. The final substitution priority $P_{final}$ is calculated by applying the modifier to a baseline inverse of the cumulative performance, scaled by a factor $\alpha$:
$\text{Baseline}=100\times(1.0-P_{\text{cum}})$ (1)
$P_{\text{final}}=\text{clip}\left(\text{Baseline}+(\text{Modifier}\times\alpha),\ 0,\ 100\right)$ (2)
Where $\alpha=0.25$ in the current version. This ensures that the fuzzy logic acts as a ”tuner” that intensifies or attenuates the urgency based on tactical context (e.g., a yellow card or a goal scored), rather than overriding the player’s core performance metric entirely. This methodology allows the expert’s knowledge ESPN_s_d; Terra_s_d; CNN_s_d to be faithfully captured by the system, adding significant weight to the player’s performance on the field.
4.7 A Normative Taxonomy of Action-Oriented AI Systems: Formalizing Prescriptive AI
The ability to recommend or evaluate actions has been addressed across multiple research traditions and industrial domains under diverse labels, including decision optimization, action recommendation, prescriptive analytics, and AI-driven decision support. These approaches are often discussed as if they constituted a coherent class of systems. In reality, they occupy a heterogeneous and only loosely structured conceptual space, within which crucial distinctions regarding uncertainty, accountability, and human agency are frequently overlooked.
This section establishes a coherent conceptual hierarchy for action-oriented decision systems. We first characterize the broad and heterogeneous superset of action-oriented systems. We then situate Prescriptive Analytics as a structured decision-oriented framework within that space. Finally, we formally delineate Prescriptive Artificial Intelligence (Prescriptive AI) as a normative, high-stakes subset of Prescriptive Analytics, defined not by algorithmic choices but by its epistemological role in human decision-making.
4.7.1 Action-Oriented Decision Systems: A Heterogeneous Superset
Action-oriented decision systems form a broad superset of computational approaches whose shared property is the capacity to recommend, prioritize, or execute actions. This superset includes, among others:
  • Optimization and mathematical programming systems;
  • Simulation-based and what-if analysis tools;
  • Rule-based and expert systems;
  • Reinforcement Learning (RL) agents;
  • Recommendation and next-best-action systems;
  • AI-augmented decision support tools embedding predictive models into workflows.
Despite their shared action-producing nature, these systems differ substantially in objectives, assumptions, and interaction with human decision-makers. Many are designed primarily for automation or outcome maximization, often assuming stable objectives, well-defined environments, and limited requirements for interpretability or accountability. Consequently, action-oriented systems do not constitute a single paradigm, but rather a conceptually heterogeneous collection of approaches.
4.7.2 Prescriptive Analytics as a Decision-Oriented Subset
Within this heterogeneous landscape, Prescriptive Analytics provides a more structured and widely recognized framework. Originating in business intelligence and operations research, prescriptive analytics focuses on determining what actions should be taken to achieve specified objectives, typically through optimization, simulation, and rule-based reasoning Davenport2007; Lepenioti2020; Bertsimas2020.
Prescriptive Analytics narrows the scope of action-oriented systems by explicitly addressing decision recommendation. However, it remains largely agnostic with respect to how uncertainty, human judgment, or accountability should be handled. It encompasses both automated and decision-support systems and does not inherently distinguish between optimizing outcomes and supporting human deliberation.
Accordingly, Prescriptive Analytics is best understood as a decision-oriented subset of action-oriented systems: analytically grounded, but normatively under-specified.
Action-Oriented AI Systems (The Heterogeneous Superset) Reinforcement Learning (Automation-Oriented) Recommender Systems Predictive Workflows Prescriptive Analytics (Decision-Oriented) Black-Box Optimization Simulation-Based Planning Prescriptive AIDecision AuditingIntrinsic InterpretabilityRisk-Aware Reasoning
Figure 5: A normative taxonomy of action-oriented AI systems. The diagram illustrates the hierarchical relationship where Action-Oriented Systems form the superset. Prescriptive Analytics sits within as a decision-focused subset, while Prescriptive AI serves as the normative core, defined by auditability and interpretability requirements.
4.7.3 Formal Definition of Prescriptive AI
Within Prescriptive Analytics, we define Prescriptive Artificial Intelligence (Prescriptive AI) as a specialized normative subset designed explicitly for high-stakes, human-in-the-loop decision-making.
Definition. Prescriptive AI is a class of intelligent systems whose primary function is to audit, justify, and support human decisions under uncertainty, rather than to automate actions or predict future outcomes.
Prescriptive AI systems are not characterized by the techniques they employ, but by the role they play in the decision process. In particular, they are distinguished by the following defining properties:
  • A primary function of decision auditing rather than decision automation;
  • Explicit reasoning under uncertainty and asymmetric risk;
  • Intrinsic interpretability as a structural requirement, not a post-hoc feature Rudin2019;
  • Contestable recommendations that preserve human agency;
  • Evaluation criteria that decouple state assessment from outcome realization, explicitly avoiding outcome bias BaronHershey1988.
In its lexical sense, a “prescription” is the action of laying down authoritative rules or directions, or more generally, something prescribed as a rule mw:prescription. By this definition, a prescriptive statement is inherently normative: it asserts what ought to be done rather than merely predicting or suggesting an outcome. Consequently, any system that cannot be inspected, audited, or contested cannot properly prescribe.
Taken together, these properties establish a necessary boundary: a system that cannot be audited cannot prescribe; it can only suggest or automate. These properties impose normative constraints that exclude many systems commonly labeled as prescriptive. Prescriptive AI therefore constitutes a proper subset of Prescriptive Analytics, defined by epistemological and accountability requirements rather than by optimization or learning paradigms.
4.7.4 Decision-Making as Epistemic State Transition
The normative necessity of Prescriptive AI follows from a dynamic conception of rationality. In the framework of logical dynamics developed by van Benthem, decision-making is not primarily evaluated by realized outcomes, but by the epistemic state transitions induced by actions under partial and evolving information (vanBenthem2007; vanBenthem2011). Rationality, under this view, is a property of how an agent updates beliefs, constraints, and commitments at the moment of action, rather than of the stochastic realization of downstream consequences.
This perspective provides a formal grounding for the core design principles of Prescriptive AI. If decision quality is defined at time $t$, prior to outcome realization at time $t+n$, then systems designed to support or evaluate human decisions must operate on the informational state available at the moment of commitment. Predictive accuracy and ex post outcome optimization are therefore insufficient as normative criteria. A prescriptive system must instead audit whether a decision constitutes a coherent, risk-aware, and justifiable epistemic update given the available evidence and constraints.
From this standpoint, Prescriptive AI can be understood as an operational instantiation of epistemic action auditing. Rather than attempting to predict or automate behavior, the system evaluates the internal consistency of the decision state itself, including the alignment between observed signals, inferred risk, and permissible actions. This aligns directly with the logical-dynamic view in which actions are epistemic interventions that transform an agent’s informational state.
While the present work does not implement the full formal machinery of Dynamic Epistemic Logic, it adopts this foundational stance by treating decision evaluation as a state-based rather than outcome-based problem. The logical dynamics literature does not prescribe a specific algorithmic realization; instead, it characterizes the object of evaluation. Prescriptive AI derives its epistemic effectiveness precisely from this characterization: by auditing epistemic transitions rather than stochastic outcomes, it enables principled assessment of decision quality under uncertainty, irreversibility, and asymmetric risk.
4.7.5 Prescriptive AI vs. Predictive and Outcome-Driven Systems
Predictive AI systems, typically trained via supervised learning, map contextual inputs to historical human decisions or observed outcomes. While effective for forecasting and pattern discovery, such models inherently learn and reproduce historical behavior, including embedded cognitive biases such as status quo bias, sunk cost fallacy, and outcome bias.
Prescriptive AI systems, by contrast, operate as decision auditors. Rather than imitating past decisions, they evaluate the current state against explicit objectives, constraints, and domain knowledge. Their purpose is to assess whether a decision is justified given the information available at time $t$, not whether it coincidentally leads to a favorable outcome at time $t+n$.
This distinction is critical in stochastic, high-stakes environments. Evaluating decision systems based on realized outcomes introduces outcome bias. A high-risk decision state is not retroactively validated by a favorable random outcome. Prescriptive AI explicitly decouples state evaluation from outcome realization, treating disagreement with historical decisions or subsequent outcomes not as error, but as potential identification of latent risk.
Dimension Predictive AI Prescriptive Analytics Prescriptive AI
Primary Objective Forecast events or actions Recommend actions to optimize objectives Support normative human decision-making
Learning Paradigm Supervised / statistical learning from historical data Optimization, simulation, rules, or learned models Rule-based, symbolic, or hybrid (learning-agnostic)
Optimization Target Accuracy / loss minimization Expected utility or objective maximization Risk mitigation / decision justification
Relation to Bias Learns and reproduces historical bias Implicitly reflects modeled objectives Audits, constrains, and exposes bias
Evaluation Criterion Predictive accuracy Outcome-based performance State-based validity / risk detection
Outcome Dependence Strong (labels and metrics) Strong or implicit Explicitly decoupled from outcomes
Interpretability Optional / post-hoc Often secondary Structural requirement
Role of Human Source of labels Executor or overseer Accountable decision-maker
Table 7: Conceptual comparison between predictive AI, prescriptive analytics, and prescriptive AI paradigms. Prescriptive AI is distinguished not by a specific learning paradigm, but by its normative role in auditing and supporting human decisions under uncertainty.
4.7.6 Boundary Cases Within Prescriptive Analytics
Several classes of systems fall within Prescriptive Analytics but do not satisfy the defining criteria of Prescriptive AI:
Optimization-based systems compute actions that maximize objective functions under constraints Bertsimas2020. While effective in structured settings, they typically assume stable objectives and do not audit human decision justification under uncertainty.
Reinforcement Learning systems prescribe actions by optimizing expected reward SuttonBarto2018. Despite their prescriptive nature, they are often opaque and automation-oriented, limiting contestability and interpretability, and are particularly vulnerable to outcome bias in stochastic environments.
Recommendation and next-best-action systems prioritize actions based on engagement or conversion metrics, relying on outcome-driven evaluation and rarely addressing accountability or asymmetric risk.
Rule-based and expert systems offer interpretability but often rely on static knowledge and lack mechanisms for uncertainty-aware temporal auditing.
These systems are prescriptive in an analytic sense, but do not fulfill the normative role required of Prescriptive AI.
4.8 Hybrid Systems and Overlapping Set Membership
The conceptual sets described here are not mutually exclusive. Hybridization is the norm in modern AI systems, which frequently combine symbolic reasoning, optimization, statistical learning, and deep learning.
As with the relationship between machine learning and deep learning, subset membership reflects conceptual scope rather than exclusivity. A system may employ optimization or learning techniques associated with Prescriptive Analytics while simultaneously satisfying the normative criteria that qualify it as Prescriptive AI. Classification therefore depends not on the techniques employed, but on the system’s functional role in the decision-making process.
4.8.1 Cross-Domain Applicability: From Pitch to Critical Operations
While elite-level soccer serves as an observable laboratory for adversarial decision-making, the structural constraints of the domain—namely irreversibility, severe time pressure, and asymmetric risk—are isomorphic to those encountered in several safety-critical sectors. The prescriptive auditing framework proposed herein passos2025 does not rely on domain-specific mechanics, but on the temporal auditing of decision-relevant performance signals and agent state evolution. Consequently, the architecture is transferable to environments in which human judgment is susceptible to fatigue, status quo bias, or desensitization under sustained cognitive load.
  • Clinical Triage and High-Dependency Care: In emergency and critical-care settings, clinicians operate under persistent cognitive load, where subtle signs of deterioration may be overlooked due to desensitization or alarm fatigue. A prescriptive auditor could operate over continuous indicators of physiological state evolution—analogous to the performance signals used in this study—to surface latent risk trajectories that remain within nominal ranges but exhibit adverse momentum. Rather than producing diagnostic predictions, the system would prescribe prioritized re-evaluation, supporting timely intervention before irreversible deterioration occurs.
  • Financial Risk Management (Trading Desks): In financial decision-making, human operators are known to exhibit systematic biases such as sunk cost effects and loss aversion, leading to prolonged exposure to unfavorable positions. A prescriptive auditing system could continuously evaluate the alignment between risk exposure and market dynamics, issuing normative recommendations when exposure exceeds logic-based thresholds. In this sense, the system functions not as a predictor of market outcomes, but as an auditor of decision states under volatility.
  • Aviation and Control Room Operations: In air traffic control and other safety-critical control rooms, operator fatigue and cognitive degradation constitute latent failure modes that are difficult to observe directly. A prescriptive auditor could assess indicators of cognitive performance degradation—such as response latency or interaction patterns—to recommend preemptive role rotation or task redistribution before safety margins are compromised. The “substitution” logic instantiated in this work thus generalizes to proactive management of human operational capacity.
Across these domains, the core function of the system remains invariant: to decouple the evaluation of the decision state at time $t$ from stochastic outcomes at time $t+n$, thereby mitigating outcome bias in high-variance, high-stakes environments.
4.8.2 Terminological Convergence and Alternative Taxonomies
Current developments in machine intelligence exhibit a structural pivot from purely predictive architectures toward systems designed to constrain, guide, or audit agentic behavior. This convergence reflects the growing recognition that stochastic outcome prediction is frequently insufficient for evaluating performance in high-stakes, dynamic environments where process validity and contextual coherence supersede result optimization. We introduce the term Prescriptive AI to describe this emergent class of systems. Unlike predictive models that estimate the probability of a future state, Prescriptive AI establishes a normative framework for evaluating the coherence of a decision at the precise moment of its execution.
This perspective aligns with foundational work on the logical dynamics of information and agency, particularly van Benthem’s account of decision-making as an epistemic state transition under uncertainty (vanBenthem2007; vanBenthem2011). While the present work does not implement the formal machinery of dynamic epistemic logic, it adopts this conceptual stance by evaluating decisions based on the informational state and normative criteria available at the moment of action, rather than on the stochastic realization of downstream outcomes. By treating decision-making as a dynamic process of information uptake and evaluation, Prescriptive AI analytically decouples the agent’s reasoning from stochastic environmental variability, enabling principled auditing of decision quality independently of realized results.
This definition allows us to position Prescriptive AI precisely relative to adjacent taxonomies. While we adopt this term to emphasize the normative function of the system, the framework intersects with, and provides empirical grounding for, several established conceptual traditions:
  • Decision Intelligence (DI): Commonly used in industrial and organizational contexts, Decision Intelligence emphasizes the engineering of decision workflows. The proposed framework constitutes a specific instantiation of DI in which the critical decision interface is mediated by intrinsically interpretable symbolic reasoning, addressing auditability requirements that are often under-specified in conventional DI implementations.
  • Normative AI: Within the ethics and governance literature, systems that encode explicit standards of acceptable behavior are often described as normative. The system employed in this work functions as a normative agent insofar as it enforces a codified standard of decision quality—derived from both hard performance constraints and soft contextual heuristics—against which human actions are evaluated in real time.
  • Cognitive Orthotics: In human–computer interaction, cognitive orthotics refer to systems designed to compensate for specific human cognitive limitations, such as attentional lapses or systematic biases. By explicitly correcting for effects such as status quo bias and outcome bias, the proposed framework operates as a real-time cognitive stabilization mechanism in high-stakes decision contexts.
  • Real-Time Algorithmic Auditing: Algorithmic auditing is traditionally framed as a post-hoc activity concerned with fairness or compliance. In contrast, the present approach extends auditing into the decision loop itself, enabling continuous, run-time evaluation of human decision states prior to the execution of irreversible actions.
Within this landscape, Prescriptive AI serves as a unifying descriptor that emphasizes the system’s functional output—the prescription—while retaining the requirement for intrinsic interpretability and auditability that is central across these literatures.
4.8.3 Implications for High-Stakes Decision Making
By explicitly positioning Prescriptive AI as a normative subset of Prescriptive Analytics, and Prescriptive Analytics as a structured subset of a broader heterogeneous space of action-oriented systems, this work resolves a persistent source of conceptual confusion in both academic and industrial discourse Sun2020; Wissuchek2022.
This hierarchy clarifies that:
  • Not all action-oriented systems are prescriptive analytics;
  • Not all prescriptive analytics systems are Prescriptive AI;
  • Hybrid systems may legitimately belong to multiple conceptual sets.
Prescriptive AI should therefore be understood not as a generic label for action recommendation, but as a distinct AI paradigm designed to audit and support human judgment under uncertainty. In the remainder of this work, elite soccer is employed not merely as a case study, but as a demanding stress-test environment to operationalize and validate these principles under time pressure, incomplete observability, and asymmetric risk.
4.9 Ethical Use of AI in High-Stakes Decision Systems
Auditing artificial intelligence systems in high-stakes decision-making contexts is not merely a technical challenge but a fundamentally ethical one. As argued by Rudin2019, deploying opaque, black-box models in domains where decisions have irreversible consequences is inherently problematic. When such systems are extended to environments with intense public scrutiny and continuous human judgment—such as elite professional soccer—the ethical risks are amplified rather than mitigated.
In contrast to low-stakes or background automation, in-game tactical decisions are subject to immediate visual auditing by multiple human observers, including coaches, players, analysts, referees, and millions of spectators. Any divergence between an algorithmic recommendation and human intuition is instantly questioned. In this setting, the adoption of black-box models poses a substantial risk of rejection, driven by two central ethical concerns: epistemic opacity and institutional distrust.
Epistemic opacity and public accountability.
The vast majority of stakeholders involved in professional sports—including coaches, athletes, and fans—do not possess formal training in artificial intelligence, let alone in complex models such as deep neural networks. When a black-box system produces a recommendation that contradicts human judgment, the absence of a transparent rationale forces stakeholders to accept the decision on faith alone. This requirement for blind trust is ethically untenable in high-stakes environments, where decisions can alter careers, financial outcomes, and competitive integrity.
In contrast, auditable and intrinsically interpretable models—such as fuzzy logic systems—enable post-hoc and real-time explanation of decisions at multiple levels of abstraction. In the proposed framework, the entire reasoning pipeline remains open to inspection: from the microscopic computation of performance scores, through fuzzy membership functions, to the macroscopic aggregation of linguistic rules that generate the final substitution priority. This transparency substantially reduces epistemic asymmetry between the system and its users, fostering informed trust rather than blind acceptance. Importantly, such auditability allows non-technical stakeholders to understand why a recommendation was made, even without understanding *how* the underlying mathematics operates.
Institutional distrust and risk of misuse.
High-stakes domains with significant financial flows are particularly vulnerable to suspicion, manipulation, and perceived conflicts of interest. Professional sports exemplify this risk. The rapid expansion of online betting markets has reshaped the economic landscape of elite soccer, with betting companies now acting as major sponsors of clubs and leagues. Concurrently, multiple high-profile cases of match manipulation and betting-related misconduct involving professional players have eroded public trust.
In such an environment, the deployment of opaque AI systems introduces severe ethical hazards. Consider a scenario in which a black-box model recommends substituting a star player during a decisive match, followed by a negative outcome. Without an explicit and intelligible justification, the decision becomes vulnerable to accusations of corruption, collusion with betting interests, or deliberate sabotage. Coaches would lack defensible grounds for their actions, players could be unfairly penalized without explanation, and clubs would face reputational and legal risks. These dynamics create strong institutional incentives to avoid black-box systems altogether, regardless of their predictive accuracy.
By contrast, prescriptive and interpretable decision support systems mitigate these risks by design. Rather than replacing human judgment, they aim to *augment* it, offering structured, explainable insights that the human decision-maker may accept, reject, or contextualize. This human-in-the-loop paradigm significantly reduces the probability of catastrophic errors while preserving accountability. Moreover, because every recommendation can be traced to explicit rules and contextual indicators, the system provides a defensible audit trail that protects coaches, players, and institutions alike.
Prescriptive AI as an ethical design principle.
The ethical challenge in high-stakes AI is not merely accuracy but responsibility. Predictive models trained to imitate historical decisions inevitably inherit human biases, including status quo bias, outcome bias, and risk aversion. In contrast, prescriptive systems grounded in symbolic reasoning evaluate current conditions against normative rules, enabling systematic auditing of human decisions rather than their replication.
This work adopts interpretability and prescriptiveness as core ethical design principles. By prioritizing transparency, auditability, and human agency, the proposed system aligns with emerging arguments that intrinsically interpretable models are not only preferable but necessary in domains where errors carry disproportionate consequences (rudin). Rather than automating authority, the system enhances human decision-making, offering a practical pathway for deploying AI responsibly in high-stakes environments.
Ultimately, while the adoption of AI in decision support is unavoidable, its ethical deployment is not optional. In contexts such as elite soccer—where decisions are public, consequential, and irreversible—only transparent, auditable, and human-centered systems can achieve legitimate and sustainable integration.
5 Acknowledgments
In the spirit of Open Science, the complete source code, including pre-processing pipelines, fuzzy inference engines, and validation scripts, is publicly available at https://github.com/Pedro-Passos77/AI-Assisted-Substitution-Decisions-A-Fuzzy-Logic-Approach-to-Real-Time-Game-Management. This manuscript constitutes a substantially expanded version of the research accepted by the Wharton Sports Analytics Journal and the accepted in the AAAI 2026 Bridge on LMreasoning) Passos2025_Wharton passos2026, with improvements focused on methodological robustness, theoretical foundation, proposed paradigm and extended case study. We gratefully acknowledge Luca Pappalardo and collaborators for making the Soccer match event dataset publicly available (Pappalardo2019), which was essential for this study. Generative AI tools were used to assist in the drafting and structural organization of this manuscript under strict human supervision; the authors remain fully responsible for the final content and its normative justification.
5.1 Limitations and Future Work
While the proposed Fuzzy Control System (FCS) demonstrates substantial utility as a prescriptive auditing tool, it is essential to delineate its epistemic scope and operational boundaries. The system is not designed to perform causal intervention, predict exogenous events, or replace human judgment. Its function is strictly limited to auditing the observable decision space given the information available at time $t$, and should therefore be interpreted as a normative support mechanism rather than an autonomous decision-maker.
From a data perspective, the reliance on minutes_played as a proxy for physical fatigue assumes a quasi-linear accumulation of workload, serving as an estimation rather than a direct physiological measurement. This abstraction does not capture inter-individual physiological differences (e.g., genetics), variations in match intensity such as the volume of high-speed running, or metabolic expenditure under sustained defensive pressure. As a result, the model may underestimate fatigue in high-exposure contexts or overestimate wear in structurally low-intensity tactical systems.
At the modeling level, the fuzzy rule base is statically defined using expert heuristics. While this design choice is central to intrinsic interpretability and contestability, it may limit generalization across divergent tactical philosophies (e.g., high-pressing versus possession-oriented systems) without manual recalibration of membership functions or rule weights ($\alpha$). This reflects a broader trade-off between explainability and adaptive flexibility that is intrinsic to symbolic reasoning systems.
Importantly, the framework does not account for abrupt exogenous events such as acute injuries, referee decisions, or random disruptions, which occur independently of observable fatigue or performance decay. Such events lie outside the epistemic reach of any decision-auditing system grounded in pre-intervention signals.
Future iterations of this framework aim to address these limitations through three complementary research directions:
  1. 1.
    Integration of Biometric Telemetry: Incorporating GPS, accelerometer, and physiological signals to replace temporal proxies with metrics grounded in both internal and external workload, thereby refining fatigue estimation under heterogeneous match conditions.
  2. 2.
    Neuro-Fuzzy Adaptation (ANFIS): Developing an Adaptive Neuro-Fuzzy Inference System that enables dynamic calibration of membership functions and rule weights through supervised learning on historical expert decisions, preserving interpretability while improving contextual adaptability.
  3. 3.
    Tactical Profile Matching: Extending the framework beyond substitution priority to recommend replacement profiles based on detected tactical deficiencies, enabling decision support not only on whether to intervene, but also on how to intervene given available resources.