### v6

# Prescriptive Artificial Intelligence: A Formal Paradigm for Auditing Human Decisions Under Uncertainty

Abstract
We formalize Prescriptive Artificial Intelligence as a distinct paradigm for human–AI decision collaboration in high-stakes environments. Unlike predictive systems optimized for outcome accuracy, prescriptive systems are designed to recommend and audit human decisions under uncertainty, providing normative guidance while preserving human agency and accountability.
We introduce four domain-independent axioms characterizing prescriptive systems and prove fundamental separation results. Central among these is the Imitation Incompleteness theorem, which establishes that supervised learning from historical decisions cannot correct systematic decision biases in the absence of external normative signals. Consequently, performance in decision imitation is bounded by a structural bias term $\varepsilon_{\text{bias}}$ rather than the statistical learning rate $\mathcal{O}(1/\sqrt{n})$.
This result formalizes the empirically observed accuracy ceiling in human decision imitation tasks and provides a principled criterion for when automation should be replaced by epistemic auditing. We demonstrate the computational realizability of the framework through an interpretable fuzzy inference system, applied as a stress test in elite soccer decision-making, where it reveals systematic decision latency and risk states obscured by outcome and status quo biases.
The proposed framework establishes Prescriptive AI as a general, realizable class of decision-support systems applicable across safety-critical domains in which interpretability, contestability, and normative alignment are essential.
keywords:
Prescriptive Artificial Intelligence , Decision Auditing , Human–AI Collaboration , Epistemic State Transition , Outcome Bias , Fuzzy Inference Systems , Explainable Artificial Intelligence , Cognitive Bias , High-Stakes Decision Making , Normative Decision Theory
[…]
[inst1] organization=Institute of Computing, Universidade Federal Fluminense (UFF), addressline=Av. Gal. Milton Tavares de Souza, postcode=24210-346, city=Niterói, state=RJ, country=Brazil
1 Introduction
Decision making under uncertainty in high-stakes environments constitutes a central challenge for human judgment. Settings such as clinical triage, operational risk management, and strategic resource allocation are characterized by irreversible actions, asymmetric consequences, and a critical dependence on timely and well-calibrated decisions. In such domains, decision errors are costly not only because of incorrect outcomes, but because they are often unjustified, delayed, or poorly aligned with underlying risk.
Despite these constraints, computational decision-support systems have been dominated by a fundamentally predictive paradigm. Models are typically optimized to replicate historical decision patterns, achieving statistical accuracy by mimicking prior human behavior. However, reliance on behavioral imitation is problematic: historical decisions often encode cognitive biases and institutional inertia. As a result, predictive systems create a structural mismatch between accuracy and decision quality, reaching a documented “predictive ceiling”—for instance, state-of-the-art substitution models plateau at approximately $70\%$ accuracy by merely cloning human choices [26]. Moreover, standard evaluation practices in stochastic environments suffer from outcome bias, judging decision quality by realized results rather than by the epistemic justification available at the time of choice [4].
Importantly, this work does not aim to improve predictive accuracy, automate decision-making, or replace human expertise. Instead, it focuses on auditing the quality of human judgment under uncertainty by separating epistemic justification from stochastic outcomes. From this perspective, the limitations of predictive systems are not merely due to data insufficiency, but reflect a normative flaw inherent to supervised learning in agentic contexts. We argue that this limitation is fundamental: in the absence of external normative signals, decision imitation cannot, even asymptotically, correct systematic human bias. By mapping contextual inputs to historical decisions, such systems inevitably reproduce embedded cognitive biases, including status quo bias (reluctance to change strategy) and sunk cost fallacy (persistence in failing courses of action). Consequently, predictive models tend to validate conservative human behavior rather than reveal when deviation is epistemically warranted.
To address this limitation, we distinguish between Predictive AI, which forecasts events or replicates past decisions, and Prescriptive AI, defined here as a class of systems whose primary function is to audit, justify, and support human decision-making under uncertainty. This perspective aligns with decision-theoretic and epistemic accounts of agency, in which rational choice is defined as a coherent state transition at time $t$, independent of the stochastic outcome realized at time $t+n$ [48]. While prescriptive analytics and interpretable decision-support frameworks have been advocated in theory [6, 21, 39], empirical demonstrations in realistic, adversarial environments remain scarce. Black-box models are particularly ill-suited for such auditing tasks due to automation bias and opacity [32], motivating the need for systems that explicitly map observed states to epistemically justifiable actions.
In this work, we use elite-level soccer as a demanding natural laboratory to operationalize the prescriptive auditing paradigm. Substitution decisions share structural properties with many safety-critical contexts: they are time-sensitive, irreversible, and made under pervasive uncertainty. Existing approaches in this domain are limited either by insufficient temporal resolution—introducing exposure bias by favoring cumulative playing time over efficiency [30, 42]—or by a reliance on purely predictive modeling that fails to surface normative tactical risk [50, 28].
To enable systematic auditing of human judgment, we propose a hybrid statistical–symbolic framework that overlays intrinsically interpretable fuzzy reasoning onto robust statistical signals. Rather than automating decisions, the system functions as an auditing layer, surfacing decision-relevant risk—such as performance decay or defensive liability—before it manifests as observable failure. Unlike descriptive fuzzy models [23, 52] or static ranking systems [40], the proposed approach operationalizes continuous, role-aware evaluation of evolving decision states.
This work makes three contributions. First, we formalize a prescriptive auditing framework for evaluating human decision-making under uncertainty, grounded in domain-independent axioms and separation results that distinguish auditing from prediction. Second, we introduce a role-aware cumulative mean metric that eliminates play-time exposure bias, enabling principled detection of intra-episode performance deterioration. Third, we demonstrate the computational realizability of the proposed prescriptive framework in a high-stakes adversarial environment, showing that it systematically reveals latent risk patterns overlooked by both human experts and black-box predictive models.
2 Research Context and Review
This section situates the present work within the broader literature on decision theory, prescriptive analytics, and human-centered decision support. Beyond surveying domain-specific applications in sports analytics, the review deliberately incorporates foundational results from decision theory and cognitive science that formalize how decisions should be evaluated under uncertainty and how humans systematically deviate from normative rationality. These theoretical contributions provide the normative and cognitive grounding for interpreting Prescriptive AI not merely as an optimization technology, but as an auditing mechanism designed to compensate for structural limitations in human real-time decision-making.
2.1 Scope Summary of Reviewed Works
Category Representative Works
Theoretical Foundations
Prescriptive Analytics & Optimization [21]; [8]; [6]; [49]
Decision Support & Agency [36]; [29]; [44]; [4]; [32]
Logic of Agency & Information Dynamics [47]; [48]
Normative Decision Theory & Bounded Rationality [41]; [43]; [18]; [15]; [33]; [10]
Domain Applications
Contextual Tactical Reasoning [52]; [46]; [23]
Fuzzy Individual Evaluation [51]; [5]; [40]; [27]; [20]
Substitution Analysis & ML [16]; [37]; [28]; [50]; [26]
Performance Metrics [30]; [42]
Methodological Pillars
Explainable AI (XAI) [39]; [25]; [9]; [22]; [17]; [19]
Table 1: Scope-oriented categorization of related work, spanning theoretical foundations, domain-specific applications, and methodological pillars.
2.2 Theoretical Framework: Prescriptive AI and Decision Agency
To rigorously position the proposed system within the existing literature, it is necessary to distinguish between predictive and prescriptive paradigms of decision support. Prescriptive Analytics is commonly described as the highest stage of analytics maturity, addressing the question “what should be done?” through optimization, simulation, and rule-based reasoning [8, 21]. Recent systematic reviews confirm that while predictive models estimate future probabilities, prescriptive systems explicitly map observed states to recommended actions [49, 6], evolving historically from static Decision Support Systems (DSS) toward more dynamic and intelligent decision agents [29].
Despite this evolution, a critical gap persists in how prescriptive systems are evaluated in stochastic environments. Traditional approaches, including standard reinforcement learning (RL) agents [45], often exhibit outcome bias, in which the quality of a decision is judged primarily by its realized result rather than by the reasoning process that produced it [4]. In high-stakes domains involving human decision-makers, such retrospective evaluation is frequently insufficient. Moreover, the well-documented risk of automation bias, whereby users over-rely on opaque algorithmic recommendations [32], further motivates the need for prescriptive systems that function as interpretable decision auditors rather than black-box oracles.
At a conceptual level, these concerns resonate with work in the logical dynamics of information and agency, particularly van Benthem’s account of decision-making as an epistemic state transition under uncertainty [47, 48]. Within this tradition, rationality is not defined by the optimality of outcomes, but by the coherence of an action with respect to the informational state and constraints available at the moment the decision is taken. Decisions are thus evaluated ex ante, independently of the stochastic realization of downstream consequences.
This perspective provides a theoretical basis for separating decision quality at time $t$ from observed outcomes at time $t+n$, a distinction that has been increasingly explored across research communities concerned with decision agency, interpretability, and accountability. Rather than focusing exclusively on outcome prediction, these lines of work emphasize the importance of assessing whether decisions are contextually justified given the available evidence and risk structure, particularly in environments characterized by uncertainty, irreversibility, and delayed feedback.
From a normative standpoint, this ex-ante evaluation of decision quality is rooted in classical decision theory. Savage’s axiomatization of rational choice formalizes decision optimality as a function of expected utility conditional on the informational state available at time $t$, independently of future realizations [41]. From a normative standpoint, this ex-ante evaluation of decision quality is rooted in classical decision theory. Savage’s axiomatization of rational choice formalizes decision optimality as a function of expected utility conditional on the informational state available at time $t$, independently of future realizations [41]. In this sense, van Benthem defines the epistemic state underlying a decision, while Savage provides the normative evaluation over that state. Complementarily, the theory of bounded rationality introduced by Simon [43] demonstrates that human decision-makers are structurally incapable of performing such continuous optimization under time pressure. Empirical models of belief updating further show that sequential evidence integration is systematically biased by order and salience effects [18], while fast-and-frugal heuristics dominate real-time human judgment under uncertainty [15]. Together, these results reinforce the need for prescriptive systems that evaluate decision states normatively while preserving human agency, in line with established taxonomies of human–automation interaction [33] and theories of situation awareness [10].
2.3 Contextual Interpretation and Tactical Reasoning
Several studies model contextual indicators to support tactical interpretation in sports. [52] employ fuzzy contextual reasoning to produce action-oriented decisions, mapping continuous match descriptors into symbolic control outputs. Similarly, [23] applies fuzzy control systems to tactical arbitration under uncertainty. In contrast, [46] focus on the semantic recognition of tactical actions using fuzzy models, remaining descriptive rather than prescriptive. While these contributions advance contextual interpretation, they neither integrate individual performance evaluation nor address substitution decisions.
2.4 Fuzzy-Based Individual Performance Evaluation
A complementary line of research applies fuzzy logic to individual player evaluation. [51] and [5] employ fuzzy inference systems to assess player suitability and generate rankings. Subsequent works, such as [40] and [27], extend this paradigm through fuzzy and neuro-fuzzy models to synthesize performance-related variables. [20] further demonstrate the suitability of fuzzy logic for transforming physical indicators into interpretable assessments. These approaches provide interpretability but generally lack temporal dynamics or prescriptive substitution reasoning.
2.5 Substitution Analysis: From Prediction to Prescription
Research on player substitutions has traditionally followed two paths: observational analysis and predictive modeling. Studies such as [16], [37], [28], and [50] analyze substitutions using statistical, observational, or causal methods to understand timing and patterns. While valuable for post-match analysis, they do not directly support real-time decision-making.
More recently, [26] framed substitutions as a supervised prediction problem, applying machine learning models to estimate when a substitution is likely to occur. Despite achieving predictive accuracy of approximately 70%, this line of work emphasizes imitation of historical decisions rather than prescriptive optimization. Unlike Reinforcement Learning approaches that maximize a reward signal [45], which can be noisy in low-scoring sports, our approach focuses on the normative evaluation of the need for substitution, independent of the match’s stochastic outcome.
2.6 Performance Evaluation and Temporal Exposure Bias
A central challenge in multi-agent environments is quantifying individual contributions. The PlayeRank framework [30] defines a multidimensional metric validated against professional scouts. Although effective, its cumulative formulation reintroduces exposure bias, preventing the detection of performance decay or momentum reversals. Recent work by [42] demonstrates that player influence fluctuates meaningfully during a match due to tactical factors. Building on this, the present work adopts a role-aware cumulative mean over fixed temporal slices, addressing the exposure bias inherent in cumulative sums.
2.7 Symbolic Reasoning and Explainable AI (XAI)
The interpretation of imprecise information is critical for accountability. While “black-box” models dominate distinctive tasks, they are often unsuitable for high-stakes decision auditing [39]. Foundational works by [9], [22], and [25] emphasize the distinction between interpretability and post-hoc explanation. Surveys by [17] and [19] reinforce the need for transparency in human-centric AI. [39] specifically warns against explaining black boxes in high-stakes contexts, advocating for models where reasoning is transparent by design. This principle motivates our use of Fuzzy Logic not just as a controller, but as a semantic layer that ensures the system’s recommendations are contestable and auditible by the human coach.
2.8 Identified Research Gap
Despite substantial progress in both general analytics and sports science, the literature reveals three critical, intersecting gaps that this work addresses.
1. Lack of Formal Definition for Prescriptive AI: First and foremost, there is a lack of a formalized, normative definition for “Prescriptive AI.” While the term appears in recent industrial reports and applied solutions (e.g., [44]), the academic literature often conflates prescription with automation or simple prediction. There is virtually no framework that defines Prescriptive AI not just as a technological stack, but as a normative agent responsible for auditing the decision logic itself [21, 49].
2. Methodological Limitations in Dynamic Environments: Standard approaches to generating actionable insights, such as Reinforcement Learning, struggle with outcome bias [4]. In low-scoring, high-variance domains like soccer, a “good” decision can lead to a bad result (and vice-versa). Existing systems that rely on maximizing reward signals [45] or mimicking historical substitutions [26] fail to decouple the quality of the decision state from the stochastic outcome, rendering them unreliable for objective auditing.
3. Absence of Interpretable Decision Support in Soccer: Finally, the domain-specific literature remains fragmented. Soccer analytics focuses heavily on descriptive metrics [30] or predictive modelling, neglecting the prescriptive auditing of tactical decisions. No prior work jointly integrates (i) a temporally resolved, exposure-aware performance metric, (ii) intrinsic interpretability that mitigates automation bias [32], and (iii) a fuzzy reasoning engine capable of handling qualitative context.
The present work addresses these gaps by unifying temporal evaluation with symbolic decision modeling, effectively proposing a formal structure for Prescriptive AI: a system that evaluates decision quality via interpretable norms, independent of the stochastic match outcome.
3 A Normative Taxonomy of Action-Oriented Decision Systems: Formalizing Prescriptive AI
The ability to recommend or evaluate actions has been addressed across multiple research traditions and industrial domains under diverse labels, including decision optimization, action recommendation, prescriptive analytics, and AI-driven decision support. These approaches are often discussed as if they constituted a coherent class of systems. In reality, they occupy a heterogeneous and only loosely structured conceptual space, within which crucial distinctions regarding uncertainty, accountability, and human agency are frequently overlooked.
This section establishes a coherent conceptual hierarchy for action-oriented decision systems. We first characterize the broad and heterogeneous superset of Action-Oriented Decision Systems. We then situate Prescriptive Analytics as a structured decision-oriented framework within that space. Finally, we formally delineate Prescriptive Artificial Intelligence (Prescriptive AI) as a normative, high-stakes subset of Prescriptive Analytics, defined not by algorithmic choices but by its epistemological role in human decision-making.
Action-Oriented Decision Systems (The Heterogeneous Superset) Reinforcement Learning (Automation-Oriented) Recommender Systems Predictive Workflows Prescriptive Analytics (Decision-Oriented) Black-Box Optimization Simulation-Based Planning Prescriptive AIDecision AuditingIntrinsic InterpretabilityRisk-Aware Reasoning
Figure 1: A normative taxonomy of Action-Oriented Decision Systems The diagram illustrates the hierarchical relationship where Action-Oriented Decision Systems form the superset. Prescriptive Analytics sits within as a decision-focused subset, while Prescriptive AI serves as the normative core, defined by auditability and interpretability requirements.
3.0.1 Action-Oriented Decision Systems: A Heterogeneous Superset
Action-Oriented Decision Systems form a broad superset of computational approaches whose shared property is the capacity to recommend, prioritize, or execute actions. This superset includes, among others:
  • 1.
    Optimization and mathematical programming systems;
  • 2.
    Simulation-based and what-if analysis tools;
  • 3.
    Rule-based and expert systems;
  • 4.
    Reinforcement Learning (RL) agents;
  • 5.
    Recommendation and next-best-action systems;
  • 6.
    AI-augmented decision support tools embedding predictive models into workflows.
Despite their shared action-producing nature, these systems differ substantially in objectives, assumptions, and interaction with human decision-makers. Many are designed primarily for automation or outcome maximization, often assuming stable objectives, well-defined environments, and limited requirements for interpretability or accountability. Consequently, action-oriented systems do not constitute a single paradigm, but rather a conceptually heterogeneous collection of approaches.
3.0.2 Prescriptive Analytics as a Decision-Oriented Subset
Within this heterogeneous landscape, Prescriptive Analytics provides a more structured and widely recognized framework. Originating in business intelligence and operations research, prescriptive analytics focuses on determining what actions should be taken to achieve specified objectives, typically through optimization, simulation, and rule-based reasoning [8, 21, 6].
Prescriptive Analytics narrows the scope of Action-Oriented Decision Systems by explicitly addressing decision recommendation. However, it remains largely agnostic with respect to how uncertainty, human judgment, or accountability should be handled. It encompasses both automated and decision-support systems and does not inherently distinguish between optimizing outcomes and supporting human deliberation.
Accordingly, Prescriptive Analytics is best understood as a decision-oriented subset of Action-Oriented Decision Systems: analytically grounded, but normatively under-specified.
3.0.3 Formal Definition of Prescriptive AI
Within Prescriptive Analytics, we define Prescriptive Artificial Intelligence (Prescriptive AI) as a specialized normative subset designed explicitly for high-stakes, human-in-the-loop decision-making.
Definition
Prescriptive AI is a subclass of what is traditionally referred to as prescriptive analytics within the broader class of Action-Oriented Decision Systems, formally characterized by a normative decision operator whose role is to determine optimal actions and audit decision coherence under uncertainty, rather than to automate decisions or predict outcomes. This operator admits two equivalent temporal instantiations.
Norm-Invariant Recommendation and Auditing
Prescriptive AI employs a single normative decision framework for both recommendation and auditing. In prospective use, the system operates online, producing action recommendations from the epistemic state available at decision time. In retrospective analysis, the same framework is applied offline to historical records in a backtesting setting.
Auditing consists of reconstructing, from logged data, the information that was available at the moment a decision was made and recomputing the recommendation that would have been generated under the same normative criteria. Future observations are used exclusively to denoise or complete past measurements (e.g., correcting sensor errors or filling missing values), and never to introduce outcome knowledge.
Decision coherence is then assessed by comparing the reconstructed normative recommendation with the action actually taken. No separate temporal operator or learning procedure is introduced: auditing is simply norm-invariant backtesting. The same prescriptive logic governs both real-time recommendation and retrospective evaluation; only the data access mode differs.
Prescriptive AI systems are not characterized by the techniques they employ, but by the normative role they play in the decision process. In particular, they are distinguished by the following defining properties:
  • 1.
    A primary function of decision recommendation or decision recommendation rather than decision automation;
  • 2.
    Explicit reasoning under uncertainty and asymmetric risk;
  • 3.
    Intrinsic interpretability as a structural requirement, not a post-hoc feature [39];
  • 4.
    Contestable recommendations that preserve human agency;
  • 5.
    Evaluation criteria that decouple state assessment from outcome realization, explicitly avoiding outcome bias [4].
In its lexical sense, a “prescription” is the action of laying down authoritative rules or directions, or more generally, something prescribed as a rule [24]. By this definition, a prescriptive statement is inherently normative: it asserts what ought to be done rather than merely predicting or suggesting an outcome. Consequently, any system that cannot be inspected, audited, or contested cannot properly prescribe.
Taken together, these properties establish a necessary boundary: a system that cannot be audited cannot prescribe; it can only suggest or automate. These properties impose normative constraints that exclude many systems commonly labeled as prescriptive. Prescriptive AI therefore constitutes a proper subset of Prescriptive Analytics, defined by epistemological and accountability requirements rather than by optimization or learning paradigms.
3.0.4 Decision-Making as Epistemic State Transition
The normative necessity of Prescriptive AI follows from a dynamic conception of rationality. In the framework of logical dynamics developed by van Benthem, decision-making is not primarily evaluated by realized outcomes, but by the epistemic state transitions induced by actions under partial and evolving information [47, 48]. Rationality, under this view, is a property of how an agent updates beliefs, constraints, and commitments at the moment of action, rather than of the stochastic realization of downstream consequences.
This perspective provides a formal grounding for the core design principles of Prescriptive AI. If decision quality is defined at time $t$, prior to outcome realization at time $t+n$, then systems designed to support or evaluate human decisions must operate on the informational state available at the moment of commitment. Predictive accuracy and ex post outcome optimization are therefore insufficient as normative criteria. A prescriptive system must instead audit whether a decision constitutes a coherent, risk-aware, and justifiable epistemic update given the available evidence and constraints.
From this standpoint, Prescriptive AI can be understood as an operational instantiation of epistemic action auditing. Rather than attempting to predict or automate behavior, the system evaluates the internal consistency of the decision state itself, including the alignment between observed signals, inferred risk, and permissible actions. This aligns directly with the logical-dynamic view in which actions are epistemic interventions that transform an agent’s informational state.
While the present work does not implement the full formal machinery of Dynamic Epistemic Logic, it adopts this foundational stance by treating decision evaluation as a state-based rather than outcome-based problem. The logical dynamics literature does not prescribe a specific algorithmic realization; instead, it characterizes the object of evaluation. Prescriptive AI derives its epistemic effectiveness precisely from this characterization: by auditing epistemic transitions rather than stochastic outcomes, it enables principled assessment of decision quality under uncertainty, irreversibility, and asymmetric risk.
3.0.5 Prescriptive AI vs. Predictive and Outcome-Driven Systems
Predictive AI systems, typically trained via supervised learning, map contextual inputs to historical human decisions or observed outcomes. While effective for forecasting and pattern discovery, such models inherently learn and reproduce historical behavior, including embedded cognitive biases such as status quo bias, sunk cost fallacy, and outcome bias.
Prescriptive AI systems, by contrast, operate as decision auditors. Rather than imitating past decisions, they evaluate the current state against explicit objectives, constraints, and domain knowledge. Their purpose is to assess whether a decision is justified given the information available at time $t$, not whether it coincidentally leads to a favorable outcome at time $t+n$.
This distinction is critical in stochastic, high-stakes environments. Evaluating decision systems based on realized outcomes introduces outcome bias. A high-risk decision state is not retroactively validated by a favorable random outcome. Prescriptive AI explicitly decouples state evaluation from outcome realization, treating disagreement with historical decisions or subsequent outcomes not as error, but as potential identification of latent risk.
Dimension Predictive AI Prescriptive Analytics Prescriptive AI
Primary Objective Forecast events or actions Recommend actions to optimize objectives recommend normatively optimal actions and to audit decision coherence
Learning Paradigm Supervised / statistical learning from historical data Optimization, simulation, rules, or learned models Rule-based, symbolic, or hybrid (learning-agnostic)
Optimization Target Accuracy / loss minimization Expected utility or objective maximization Risk mitigation / decision justification
Relation to Bias Learns and reproduces historical bias Implicitly reflects modeled objectives Audits, constrains, and exposes bias
Evaluation Criterion Predictive accuracy Outcome-based performance State-based validity / risk detection
Outcome Dependence Strong (labels and metrics) Strong or implicit Explicitly decoupled from outcomes
Interpretability Optional / post-hoc Often secondary Structural requirement
Role of Human Source of labels Executor or overseer Accountable decision-maker
Table 2: Conceptual comparison between predictive AI, prescriptive analytics, and prescriptive AI paradigms. Prescriptive AI is distinguished not by a specific learning paradigm, but by its normative role in auditing and supporting human decisions under uncertainty.
3.1 Axiomatic Foundations of Prescriptive AI
We formalize Prescriptive Artificial Intelligence as a distinct decision-making paradigm defined by normative axioms rather than by architectures, learning procedures, or representational choices. The axioms 2 and 4 are domain-independent and specify necessary conditions that any system must satisfy in order to qualify as prescriptive, while axioms 1 and 3 operate primarily in stochastic domains
Notation.
Let $\mathcal{S}$ denote the state space and $\mathcal{A}$ the action space. Let $O_{t+n}$ denote the space of possible future outcome realizations. Let $\mathcal{N}$ denote normative criteria, including utility functions, constraints, or domain rules. Let $L$ denote a logical system used to express explicit justifications.
Definition 1 (Prescriptive System).
A prescriptive system $S$ is defined by a recommendation function
$R:\mathcal{S}\rightarrow\mathcal{A},$
and satisfies Axioms 1–4 below.
Notably, this definition does not specify: (i) how $R$ is computed (symbolic, statistical, or hybrid), (ii) the structure of $\mathcal{S}$ or $\mathcal{A}$ (discrete, continuous, or structured), (iii) whether the system learns, adapts, or remains static. Such implementation choices are orthogonal to prescriptiveness.
3.1.1 Axioms
Axiom 1 (Outcome Decoupling).
For any decision state $s_{t}\in\mathcal{S}$, the recommendation is independent of future outcome realizations:
$R(s_{t})\perp O_{t+n}.$
Note: This axiom is essential in stochastic environments where outcome realization does not reliably indicate decision quality. See Definition 1 for the distinction between outcome decoupling (prohibited) and predictive modeling of latent variables (permitted under prescriptiveness).
Axiom 2 (Epistemic Justification).
For every recommendation $a=R(s)$, there exists an explicit justification $E$, expressed in a logical system $L$, such that
$E\vdash_{L}(s\rightarrow a),$
where $L$ supports human inspection of the reasoning linking state evaluation to action selection, and $E$ is accessible to the decision-maker at decision time.
Note: The logical system $L$ may be classical propositional logic, fuzzy logic, probabilistic reasoning, structured natural language, or even implicit domain knowledge encoded during system design (e.g., literature-informed model architectures, expert-validated parameter choices). In the fuzzy implementation (Section 5.6), $L$ denotes the Mamdani inference engine, and $E$ consists of activated rules and their membership degrees; the rules themselves encode prior domain expertise from sports science literature.
Axiom 3 (State-Based Evaluation).
Decision quality is evaluated as a function of the current state and normative criteria:
$\text{Quality}(a)=f(s_{t},\mathcal{N}),$
and not as a function of realized outcomes $o_{t+n}$.
Note: In stochastic environments, this axiom ensures that decisions are evaluated by their epistemic justification at time $t$, independent of which stochastic realization occurs at time $t+n$. In the soccer implementation (Section 5.6), we instantiate $f$ as the fuzzy priority function: $\text{Quality}(a,s_{t},N):=P_{\text{final}}(a,s_{t})$ (Equation 2, Section 5.6.4).
Axiom 4 (Contestability).
Any recommendation $a=R(s)$ must satisfy:
  1. 1.
    Inspectability: the justification $E$ is accessible and comprehensible to a human agent;
  2. 2.
    Challengeability: the agent can question the validity of $E\vdash_{L}(s\rightarrow a)$;
  3. 3.
    Overridability: the agent retains authority to select an alternative action $a^{\prime}\neq a$ with documented justification.
3.1.2 General Theoretical Results
Theorem 1 (Outcome Independence).
For any prescriptive system $S$, the recommendation $R(s_{t})$ is invariant to future outcome realizations.
Proof.
The function $R$ operates exclusively on the current epistemic state $s_{t}\in\mathcal{S}$. Future outcomes belong to $O_{t+n}$ and are not elements of the domain of $R$. Therefore, for any $o_{1},o_{2}\in O_{t+n}$,
$R(s_{t}\mid o_{1})=R(s_{t}\mid o_{2}).$
$\square$
Remark (Domain of Applicability).
Outcome Independence is non-trivial only in stochastic or partially observable domains, where future outcome realizations are not deterministically entailed by the current epistemic state $s_{t}$. In fully deterministic and fully observable settings, outcome realizations collapse into state transitions and the distinction between epistemic evaluation at time $t$ and outcome observation at time $t+n$ becomes vacuous. Prescriptive AI is therefore primarily concerned with decision-making under epistemic uncertainty.
Remark (Connection to Epistemic Logic).
The axiomatic structure of Prescriptive AI aligns with van Benthem’s view of rational agency as epistemic state transition [48].
In the language of Dynamic Epistemic Logic (DEL), a prescriptive recommendation $R(s_{t})\rightarrow a$ can be interpreted as an epistemic action that transforms the agent’s information state $K_{t}$, without requiring access to future outcome realizations $O_{t+n}$.
While we do not invoke the full DEL machinery in this work, Axioms 1–3 operationalize its core principle: decision quality is a property of the epistemic transition at time $t$, rather than of stochastic realizations at time $t+n$.
Remark.
This property distinguishes prescriptive systems from outcome-based reinforcement learning, which explicitly optimizes expected returns over realized outcomes.
Theorem 2 (Imitation Incompleteness).
Systems trained solely by imitation of historical human decisions cannot, in general, correct systematic biases present in the generating policy without access to external normative or counterfactual information.
Assumptions.
  1. 1.
    The dataset $H=\{(s_{i},a_{i})\}_{i=1}^{n}$ is an IID sample from a human policy $\pi_{h}(a\mid s)$.
  2. 2.
    (Realizability) There exists a parameter $\theta^{*}$ such that $M_{\text{pred}}(\cdot;\theta^{*})=\pi_{h}(\cdot)$.
  3. 3.
    The number of samples satisfies $n\rightarrow\infty$.
  4. 4.
    $H$ contains no explicit normative or counterfactual signals.
Proof.
Suppose the human policy $\pi_{h}$ exhibits a systematic bias $\beta$ in the following sense: there exists at least one state $s\in\mathcal{S}$ and a normative utility function $U$ such that
$a^{\star}=\arg\max_{a}U(s,a),\quad a_{b}=\arg\max_{a}\pi_{h}(a\mid s),$
with $a_{b}\neq a^{\star}$ and $\pi_{h}(a_{b}\mid s)>\pi_{h}(a^{\star}\mid s)$.
Under the realizability assumption, standard consistency results for maximum likelihood estimation imply that the predictive model $M_{\text{pred}}$ converges in probability to $\pi_{h}$ as $n\rightarrow\infty$. Since $H$ contains no information about the normative utility $U$, the model has no basis for preferring $a^{\star}$ over $a_{b}$ and therefore selects
$M_{\text{pred}}(s)=a_{b}.$
Conversely, a prescriptive system optimizing $U$ computes
$S(s)=\arg\max_{a}f(s,U),$
independently of $H$. Hence, for this state $s$,
$M_{\text{pred}}(s)=a_{b}\neq a^{\star}=S(s).$
Since this argument holds for at least one state $s$, imitation learning cannot, in general, correct systematic bias. $\square$
Corollary 2.1 (Bounded Improvement via Imitation).
Let $p(s)$ denote the state distribution induced by the human policy $\pi_{h}$, and define the expected utility of a policy $\pi$ as:
$V(\pi):=\mathbb{E}_{s\sim p(s)}\left[\mathbb{E}_{a\sim\pi(\cdot|s)}[U(s,a)]\right]$
Let $\varepsilon_{\text{bias}}=U(s^{*},a^{*})-U(s^{*},a_{b})>0$ denote the utility gap for the biased state identified in Theorem 2.
Then as sample size $n\to\infty$:
$E[U(M_{\text{pred}})]\leq E[U(\pi_{h})]+o(1),$
while a prescriptive system achieves:
$E[U(\mathcal{S})]\geq E[U(\pi_{h})]+\varepsilon_{\text{bias}}.$
Proof. From Theorem 2, $M_{\text{pred}}$ converges to $\pi_{h}$ as $n\to\infty$. Since the expectation is taken over the state distribution induced by $\pi_{h}$, we have:
$\lim_{n\to\infty}E[U(M_{\text{pred}})]=E[U(\pi_{h})].$
For the prescriptive system, evaluation is computed over the normative utility function $U$ independently of historical decisions. At state $s^{*}$ where the bias occurs, $\mathcal{S}$ selects $a^{*}$ while $\pi_{h}$ selects $a_{b}$, yielding a utility gain of $\varepsilon_{\text{bias}}$. Therefore:
$E[U(\mathcal{S})]\geq E[U(\pi_{h})]+p(s^{*})\cdot\varepsilon_{\text{bias}}\geq E[U(\pi_{h})]+\varepsilon_{\text{bias}}$
where the final inequality holds since $p(s^{*})>0$ by assumption. $\Box$
This result formalizes a structural ceiling on imitation-based learning: predictive systems converge to the biased human policy in expectation, while prescriptive systems can achieve strictly higher normative utility by optimizing directly over $U$. The quantity $\varepsilon_{\text{bias}}$ represents irreducible normative loss under pure imitation, and will be shown in Theorem 5 to be fundamentally unobservable through outcome-based evaluation alone.
Theorem 3 (Necessity of Auditability).
Any non-auditable system cannot be prescriptive.
Proof.
If a system is non-auditable, it lacks an inspectable justification and therefore violates contestability (Axiom 4), contradicting the definition of prescriptiveness. $\square$
Theorem 4 (Strict but Non-Disjoint Separation of Paradigms).
Let PRED denote predictive systems and PRESC prescriptive systems. Then:
$\textsc{PRED}\cap\textsc{PRESC}\neq\emptyset,\quad\textsc{PRED}\nsubseteq\textsc{PRESC},\quad\textsc{PRESC}\nsubseteq\textsc{PRED}.$
Proof.
(i) Non-empty intersection: the intersection is non-empty by construction. Consider a Bayesian decision network that (i) models $P(o\mid s,a)$ via probabilistic inference, (ii) selects $a^{*}=\arg\max_{a}\mathbb{E}[U(a)\mid s]$ via an explicit decision rule, and (iii) provides a causal directed acyclic graph $G$ as justification, such that $E=\text{``}G\vdash_{L}(s\rightarrow a^{*})\text{''}$.
This system is predictive, as it models outcome distributions, and prescriptive, as it satisfies Axioms 1–4 via explicit state-based utility maximization with auditable causal structure.
(ii) $\textsc{PRED}\nsubseteq\textsc{PRESC}$ follows from Theorems 2 and 3.
(iii) $\textsc{PRESC}\nsubseteq\textsc{PRED}$ follows from symbolic rule-based expert systems that map states to actions without modeling outcome distributions. $\square$
Proposition 1 (Computational Realizability).
Axioms 1–4 are jointly satisfiable by multiple distinct computational architectures.
Proof (by construction).
We exhibit three structurally distinct implementations: (i) Fuzzy logic systems, where rules encode domain knowledge and activated rules constitute the justification; (ii) Bayesian decision systems, where expected utility is computed explicitly with causal graphs as justification; and (iii) constraint-based systems, where recommendations arise from explicit constraint satisfaction proofs.
All satisfy Axioms 1–4 while differing fundamentally in representation and computation. Prescriptiveness therefore characterizes a paradigm class rather than a unique architecture. $\square$
Corollary 4.1 (Black-Box Exclusion).
Black-box decision systems are not prescriptive.
Theorem 5 (Impossibility of Outcome-Based Decision Labeling in Stochastic Environments).
Let $\mathcal{S}$ be a prescriptive decision system operating in a stochastic environment—a setting where outcomes are substantially influenced by factors unknown or uncontrollable at decision time (Definition 1).
At decision time $T$, the system observes an epistemic state $E_{T}$ consisting of all information available at time $T$. Let $O_{T+N}(a)$ denote the stochastic outcome realized at future time $T+N$ after action $a$ is taken.
Assume:
  1. 1.
    Exactly one action $a^{*}\in A$ is executed at time $T$.
  2. 2.
    For any alternative action $a^{\prime}\neq a^{*}$, $O_{T+N}(a^{\prime})$ is counterfactual and unobservable.
  3. 3.
    Multiple actions may lead to acceptable or unacceptable outcomes under different stochastic realizations.
  4. 4.
    Prescriptive optimality is defined over $U(E_{T},a)$, not over realized outcomes $O_{T+N}$.
Then:
  1. 1.
    There exists no binary labeling function $y:(E\times A)\to\{0,1\}$ assigning “correct” or “incorrect” decisions consistently with $U(E_{T},a)$, where $E$ denotes the space of epistemic states.
  2. 2.
    Consequently, outcome-based classification metrics (accuracy, precision, recall, F1-score) are ill-defined for prescriptive systems.
  3. 3.
    Each prescriptive decision must instead be evaluated individually at time $T$ by its logical and normative consistency with epistemic state $E_{T}$, independently of realized outcome at $T+N$.
Proof. We prove by contradiction that no binary labeling function $y:(E\times A)\to\{0,1\}$ exists.
Assume such a function exists. Then for any epistemic state $E\in\mathcal{E}$ and action $a\in A$, $y(E,a)\in\{0,1\}$ assigns a definite correctness label.
Consider two realizations of the same decision instance $(E,a^{*})$ with identical epistemic states and actions, but different outcome realizations:
  • 1.
    Instance 1: epistemic state $E$, action $a^{*}$ executed, outcome $o_{1}$ realized.
  • 2.
    Instance 2: epistemic state $E$, action $a^{*}$ executed, outcome $o_{2}$ realized, $o_{1}\neq o_{2}$.
From Assumption (3), both $o_{1}$ and $o_{2}$ may occur under identical decisions due to the stochasticity of $O(a^{*})$. From Assumption (4), optimality is determined by $U(E,a^{*})$, not by outcomes.
If $y(E,a^{*})$ varies based on which outcome is realized, then $y$ encodes outcome bias, violating Axiom 3 (Outcome Independence).
If $y(E,a^{*})$ is constant across outcome realizations, then $y$ must evaluate decisions based on $E$ alone. However, from Assumption (4), epistemic evaluation requires assessing $U(E_{T},a)$ over the full action space $A$, while empirical observation provides only $(E,a^{*},o)$, not counterfactual $(E,a^{\prime},o^{\prime})$ for $a^{\prime}\neq a^{*}$.
Therefore, no binary function $y$ consistently assigns correctness labels without either:
  1. 1.
    violating Axiom 3 (Outcome Independence), or
  2. 2.
    requiring unobservable counterfactual information.
Hence, $y$ cannot exist as defined, and outcome-based classification metrics are ill-defined for prescriptive systems. $\Box$
Remark (Scope of the Impossibility Result).
This impossibility applies specifically to the practice of assigning binary correctness labels to individual decision instances based on realized outcomes—a conflation of decision quality with stochastic luck.
It does not prohibit:
  1. 1.
    Predicting continuous variables (e.g., asset values, risk probabilities) and validating predictions via regression metrics, which measure estimation accuracy rather than decision correctness;
  2. 2.
    Using aggregate outcome distributions for model calibration at the population level, without labeling individual decisions;
  3. 3.
    Deterministic environments where outcomes are fully determined by observable state and choice, such that outcome observation constitutes valid ground truth.
The distinction between evaluating variable predictions and labeling decision correctness is fundamental to understanding the boundary between prescriptive and purely predictive paradigms (see Definition 1 and Theorem 4).
Corollary 5.1 (Invalidity of Predictive Baselines).
Let $M_{\text{pred}}$ be a predictive model evaluated via outcome-based classification metrics. Then comparing a prescriptive system $\mathcal{S}$ against $M_{\text{pred}}$ using these metrics constitutes a category error.
Proof. From Theorem 5, outcome-based classification metrics are undefined for $\mathcal{S}$, as no labeling function $y:(E\times A)\to\{0,1\}$ exists. Therefore, $M_{\text{pred}}$ and $\mathcal{S}$ cannot be compared using a common metric space, as one admits outcome-grounded labels while the other does not. $\Box$
Remark (Practical Implications).
Theorem 5 does not prohibit empirical validation of prescriptive systems. Rather, it specifies that validation must be epistemic rather than outcome-based. Valid evaluation methods include:
  • 1.
    Epistemic coherence assessment: does the recommendation $a^{*}$ align with $U(E_{T},\cdot)$?
  • 2.
    Decision latency quantification: temporal gap between risk accumulation and human response.
  • 3.
    Counterfactual case analysis: e.g., Fagner Paradox illustrates system diagnosis post-facto.
  • 4.
    Longitudinal pattern detection: Technical Anchors vs Chronic Inefficacy.
These methods evaluate decision quality at time $T$ based on $E_{T}$, satisfying the epistemic evaluation principle.
3.1.3 Boundary Cases Within Prescriptive Analytics
Several classes of systems fall within Prescriptive Analytics but do not satisfy the defining criteria of Prescriptive AI:
Optimization-based systems compute actions that maximize objective functions under constraints [6]. While effective in structured settings, they typically assume stable objectives and do not audit human decision justification under uncertainty.
Reinforcement Learning systems prescribe actions by optimizing expected reward [45]. Despite their prescriptive nature, they are often opaque and automation-oriented, limiting contestability and interpretability, and are particularly vulnerable to outcome bias in stochastic environments.
Recommendation and next-best-action systems prioritize actions based on engagement or conversion metrics, relying on outcome-driven evaluation and rarely addressing accountability or asymmetric risk.
Rule-based and expert systems offer interpretability but often rely on static knowledge and lack mechanisms for uncertainty-aware temporal auditing.
These systems are prescriptive in an analytic sense, but do not fulfill the normative role required of Prescriptive AI.
3.2 Hybrid Systems and Overlapping Set Membership
The conceptual sets described here are not mutually exclusive. Hybridization is the norm in modern AI systems, which frequently combine symbolic reasoning, optimization, statistical learning, and deep learning.
As with the relationship between machine learning and deep learning, subset membership reflects conceptual scope rather than exclusivity. A system may employ optimization or learning techniques associated with Prescriptive Analytics while simultaneously satisfying the normative criteria that qualify it as Prescriptive AI. Classification therefore depends not on the techniques employed, but on the system’s functional role in the decision-making process.
3.2.1 Terminological Convergence and Alternative Taxonomies
Current developments in machine intelligence exhibit a structural pivot from purely predictive architectures toward systems designed to constrain, guide, or audit agentic behavior. This convergence reflects the growing recognition that stochastic outcome prediction is frequently insufficient for evaluating performance in high-stakes, dynamic environments where process validity and contextual coherence supersede result optimization. We introduce the term Prescriptive AI to describe this emergent class of systems. Unlike predictive models that estimate the probability of a future state, Prescriptive AI establishes a normative framework for evaluating the coherence of a decision at the precise moment of its execution.
This perspective aligns with foundational work on the logical dynamics of information and agency, particularly van Benthem’s account of decision-making as an epistemic state transition under uncertainty [47, 48]. While the present work does not implement the formal machinery of dynamic epistemic logic, it adopts this conceptual stance by evaluating decisions based on the informational state and normative criteria available at the moment of action, rather than on the stochastic realization of downstream outcomes. By treating decision-making as a dynamic process of information uptake and evaluation, Prescriptive AI analytically decouples the agent’s reasoning from stochastic environmental variability, enabling principled auditing of decision quality independently of realized results.
This definition allows us to position Prescriptive AI precisely relative to adjacent taxonomies. While we adopt this term to emphasize the normative function of the system, the framework intersects with, and provides empirical grounding for, several established conceptual traditions:
  • 1.
    Decision Intelligence (DI): Commonly used in industrial and organizational contexts, Decision Intelligence emphasizes the engineering of decision workflows. The proposed framework constitutes a specific instantiation of DI in which the critical decision interface is mediated by intrinsically interpretable symbolic reasoning, addressing auditability requirements that are often under-specified in conventional DI implementations.
  • 2.
    Normative AI: Within the ethics and governance literature, systems that encode explicit standards of acceptable behavior are often described as normative. The system employed in this work functions as a normative agent insofar as it enforces a codified standard of decision quality—derived from both hard performance constraints and soft contextual heuristics—against which human actions are evaluated in real time.
  • 3.
    Cognitive Orthotics: In human–computer interaction, cognitive orthotics refer to systems designed to compensate for specific human cognitive limitations, such as attentional lapses or systematic biases. By explicitly correcting for effects such as status quo bias and outcome bias, the proposed framework operates as a real-time cognitive stabilization mechanism in high-stakes decision contexts.
  • 4.
    Real-Time Algorithmic Auditing: Algorithmic auditing is traditionally framed as a post-hoc activity concerned with fairness or compliance. In contrast, the present approach extends auditing into the decision loop itself, enabling continuous, run-time evaluation of human decision states prior to the execution of irreversible actions.
Within this landscape, Prescriptive AI serves as a unifying descriptor that emphasizes the system’s functional output—the prescription—while retaining the requirement for intrinsic interpretability and auditability that is central across these literatures.
3.2.2 Implications for High-Stakes Decision Making
By explicitly positioning Prescriptive AI as a normative subset of Prescriptive Analytics, and Prescriptive Analytics as a structured subset of a broader heterogeneous space of action-oriented systems, this work resolves a persistent source of conceptual confusion in both academic and industrial discourse [44, 49].
This hierarchy clarifies that:
  • 1.
    Not all action-oriented systems are prescriptive analytics;
  • 2.
    Not all prescriptive analytics systems are Prescriptive AI;
  • 3.
    Hybrid systems may legitimately belong to multiple conceptual sets.
Prescriptive AI should therefore be understood not as a generic label for action recommendation, but as a distinct AI paradigm designed to audit and support human judgment under uncertainty. In the remainder of this work, elite soccer is employed not merely as a case study, but as a demanding stress-test environment to operationalize and validate these principles under time pressure, incomplete observability, and asymmetric risk.
3.3 Ethical Use of AI in High-Stakes Decision Systems
Auditing artificial intelligence systems in high-stakes decision-making contexts is not merely a technical challenge but a fundamentally ethical one. As argued by Rudin [39], deploying opaque, black-box models in domains where decisions have irreversible consequences is inherently problematic. When such systems are extended to environments with intense public scrutiny and continuous human judgment—such as elite professional soccer—the ethical risks are amplified rather than mitigated.
In contrast to low-stakes or background automation, in-game tactical decisions are subject to immediate visual auditing by multiple human observers, including coaches, players, analysts, referees, and millions of spectators. Any divergence between an algorithmic recommendation and human intuition is instantly questioned. In this setting, the adoption of black-box models poses a substantial risk of rejection, driven by two central ethical concerns: epistemic opacity and institutional distrust.
Epistemic opacity and public accountability.
The vast majority of stakeholders involved in professional sports—including coaches, athletes, and fans—do not possess formal training in artificial intelligence, let alone in complex models such as deep neural networks. When a black-box system produces a recommendation that contradicts human judgment, the absence of a transparent rationale forces stakeholders to accept the decision on faith alone. This requirement for blind trust is ethically untenable in high-stakes environments, where decisions can alter careers, financial outcomes, and competitive integrity.
In contrast, auditable and intrinsically interpretable models—such as fuzzy logic systems—enable post-hoc and real-time explanation of decisions at multiple levels of abstraction. In the proposed framework, the entire reasoning pipeline remains open to inspection: from the microscopic computation of performance scores, through fuzzy membership functions, to the macroscopic aggregation of linguistic rules that generate the final substitution priority. This transparency substantially reduces epistemic asymmetry between the system and its users, fostering informed trust rather than blind acceptance. Importantly, such auditability allows non-technical stakeholders to understand why a recommendation was made, even without understanding *how* the underlying mathematics operates.
Institutional distrust and risk of misuse.
High-stakes domains with significant financial flows are particularly vulnerable to suspicion, manipulation, and perceived conflicts of interest. Professional sports exemplify this risk. The rapid expansion of online betting markets has reshaped the economic landscape of elite soccer, with betting companies now acting as major sponsors of clubs and leagues. Concurrently, multiple high-profile cases of match manipulation and betting-related misconduct involving professional players have eroded public trust.
In such an environment, the deployment of opaque AI systems introduces severe ethical hazards. Consider a scenario in which a black-box model recommends substituting a star player during a decisive match, followed by a negative outcome. Without an explicit and intelligible justification, the decision becomes vulnerable to accusations of corruption, collusion with betting interests, or deliberate sabotage. Coaches would lack defensible grounds for their actions, players could be unfairly penalized without explanation, and clubs would face reputational and legal risks. These dynamics create strong institutional incentives to avoid black-box systems altogether, regardless of their predictive accuracy.
By contrast, prescriptive and interpretable decision support systems mitigate these risks by design. Rather than replacing human judgment, they aim to *augment* it, offering structured, explainable insights that the human decision-maker may accept, reject, or contextualize. This human-in-the-loop paradigm significantly reduces the probability of catastrophic errors while preserving accountability. Moreover, because every recommendation can be traced to explicit rules and contextual indicators, the system provides a defensible audit trail that protects coaches, players, and institutions alike.
Prescriptive AI as an ethical design principle.
The ethical challenge in high-stakes AI is not merely accuracy but responsibility. Predictive models trained to imitate historical decisions inevitably inherit human biases, including status quo bias, outcome bias, and risk aversion. In contrast, prescriptive systems grounded in symbolic reasoning evaluate current conditions against normative rules, enabling systematic auditing of human decisions rather than their replication.
This work adopts interpretability and prescriptiveness as core ethical design principles. By prioritizing transparency, auditability, and human agency, the proposed system aligns with emerging arguments that intrinsically interpretable models are not only preferable but necessary in domains where errors carry disproportionate consequences [39]. Rather than automating authority, the system enhances human decision-making, offering a practical pathway for deploying AI responsibly in high-stakes environments.
Ultimately, while the adoption of AI in decision support is unavoidable, its ethical deployment is not optional. In contexts such as elite soccer—where decisions are public, consequential, and irreversible—only transparent, auditable, and human-centered systems can achieve legitimate and sustainable integration.
4 The General Normative Prescriptive AI Framework (GNPAF)
4.1 Logical Positioning
This section introduces the General Normative Prescriptive AI Framework (GNPAF), which specifies the minimal structural requirements that any prescriptive AI system must satisfy in order to be normatively evaluable, epistemically grounded at decision time, and independent of realized outcomes.
Within this framework, we define the General Normative Decision Auditing (GNDA) procedure, which operationalizes how concrete decisions—human or artificial—can be audited for normative coherence under uncertainty.
The distinction is intentional:
  • 1.
    GNPAF specifies what must exist structurally.
  • 2.
    GNDA specifies how auditing is performed once those structures are in place.
This separation follows directly from the axiomatic distinction between epistemic state, normative evaluation, and realized outcomes, as well as from impossibility results showing that prescriptive policies cannot, in general, be learned or validated through outcome supervision alone.
Relationship to Existing Taxonomies.
GNPAF differs fundamentally from related approaches:
  • 1.
    Decision Theory (Savage, 1954): provides axiomatic foundations for rational choice, but does not specify computational structures for real-time prescriptive auditing under bounded rationality.
  • 2.
    Prescriptive Analytics: focuses on optimizing actions with respect to objectives, but does not impose explicit epistemic state formalization or outcome decoupling.
  • 3.
    Explainable AI (XAI): adds interpretability post-hoc to predictive models, whereas GNPAF requires explainability structurally as a prerequisite for prescriptiveness.
GNPAF thus occupies a distinct position: more operational than pure decision theory, more normatively constrained than optimization-driven prescriptive analytics, and more structurally committed than post-hoc explainability frameworks.
4.2 GNPAF Structural Components
GNPAF decomposes prescriptive AI systems into six components. These components are not algorithmic steps, but logical requirements. Violating any component necessarily violates at least one axiom.
The Six GNPAF Components.
  1. 1.
    Epistemic State Specification — What is known at decision time?
  2. 2.
    Action Space Delimitation — What actions are admissible?
  3. 3.
    Normative Evaluation Mapping — How should actions be evaluated ex ante?
  4. 4.
    Decision Coherence Assessment (GNDA) — Was the chosen action normatively justified?
  5. 5.
    Explainable Justification Layer (GNDA) — Why or why not (human-contestable)?
  6. 6.
    Outcome Decoupling Validation (GNDA) — Is evaluation independent of realized outcomes?
4.2.1 Epistemic State Specification
(Axioms A1, A3; van Benthem)
A prescriptive system must explicitly define the epistemic state at the moment a decision is taken.
An epistemic state represents the totality of information legitimately available at time $t$, including:
  • 1.
    observable facts,
  • 2.
    known constraints,
  • 3.
    recognized uncertainties,
  • 4.
    institutional or procedural limitations.
It explicitly excludes:
  • 1.
    future observations,
  • 2.
    realized outcomes,
  • 3.
    information revealed after the decision.
Following van Benthem, decisions are modeled as transitions originating from epistemic states rather than as reactions to realized trajectories. Epistemic states are therefore conditions of knowledge, not feature vectors or retrospective summaries.
Without explicit epistemic state specification, ex-ante normative evaluation is ill-defined.
4.2.2 Action Space Delimitation
(Axiom A1)
Given an epistemic state $s_{t}$, the prescriptive system must characterize the set of admissible actions $\mathcal{A}(s_{t})$.
Formally:
$a\in\mathcal{A}(s_{t})\iff a\text{ is feasible, permissible, and executable given }s_{t}.$
Action spaces are inherently state-dependent:
$\mathcal{A}(s_{t})\neq\mathcal{A}(s_{t^{\prime}})\quad\text{for }t\neq t^{\prime}.$
Normative evaluation is meaningful only relative to actions that were genuinely available at decision time. Evaluating a decision against infeasible or unavailable alternatives violates ex-ante rationality.
4.2.3 Normative Evaluation Mapping
(Axiom A2; Savage)
GNPAF requires an explicit normative evaluation mapping:
$f_{\text{presc}}:(s_{t},N)\mapsto\mathbb{R}^{|\mathcal{A}(s_{t})|},$
where $N$ denotes externally specified normative criteria.
The mapping must be:
  • 1.
    conditional on the epistemic state,
  • 2.
    defined ex ante,
  • 3.
    independent of realized outcomes.
By the established impossibility results, this mapping:
  • 1.
    cannot be learned purely from outcome supervision,
  • 2.
    cannot be inferred from imitation of historical decisions,
  • 3.
    cannot be validated retrospectively by performance.
The normative mapping is not optimized, but specified. This is a defining property of prescriptive systems under GNPAF.
4.2.4 Decision Coherence Assessment
GNDA operationalizes prescriptive auditing by comparing:
  • 1.
    the action actually taken by a decision-maker,
  • 2.
    with the action(s) normatively justified in the corresponding epistemic state.
The output of auditing is structural rather than outcome-based, and includes classifications such as:
  • 1.
    coherent,
  • 2.
    delayed,
  • 3.
    partially justified,
  • 4.
    normatively inconsistent.
Because outcome-based evaluation is provably illegitimate in this setting, coherence assessment constitutes the only admissible form of prescriptive auditing.
Auditing is therefore structural, not empirical.
4.2.5 Explainable Justification Layer
(Axiom A4)
GNDA formalizes that prescriptive systems, by definition, must provide explanations that are contestable by humans and grounded in:
  • 1.
    epistemic factors,
  • 2.
    normative criteria,
  • 3.
    explicit trade-offs.
Explainability in GNPAF is not post-hoc or auxiliary. It is a structural requirement for normative accountability.
Post-Hoc XAI GNPAF
Timing After model output During decision process
Source Approximate Direct
Status Optional Axiomatic (A4)
Human role Observer Contestor
Explainability is required for accountability, not added for transparency.
4.2.6 Outcome Decoupling Validation
GNPAF requires explicit verification that none of the preceding components depends on realized outcomes or future information unknown at decision time.
Applicability: This component is essential in stochastic environments where Axiom 1 applies (see Definition 1). In deterministic settings where outcomes uniquely determine decision quality, outcome-based validation remains valid and this component may be omitted.
Validation procedure:
  1. 1.
    Identify all inputs to Components 1–5
  2. 2.
    Verify that no input depends on $O_{t+n}$ (outcomes at time $t+n$)
  3. 3.
    Confirm that any predictive models used estimate variables from state $s_{t}$, not from future realizations
This validation step prevents collapse into retrospective outcome-based evaluation and enforces the ex-ante epistemic nature of prescriptive rationality.
4.3 Graphic diagram
Figure 2: Prescriptive AI framework illustration
4.4 Status of the Framework
GNPAF is neither an optimization method nor a learning algorithm. It defines the minimal structural conditions for prescriptive AI systems under uncertainty.
GNDA provides the operational procedure for auditing decisions instantiated within that structure.
Any prescriptive system violating one GNPAF component necessarily violates at least one axiom.
4.5 Cross-Domain Applicability: From Pitch to Critical Operations
High-stakes, real-time decision-making environments provide naturalistic laboratories for observing adversarial and time-sensitive judgments. The structural constraints of such domains—namely, irreversibility of actions, severe time pressure, and asymmetric risk—are common across many safety-critical sectors. The prescriptive AI framework (GNPAF) proposed here in  [34] does not depend on domain-specific mechanics, but rather on the temporal auditing of decision-relevant performance signals and agent state evolution. By incorporating logically interpretable reasoning and principles from Explainable AI, the architecture offers a blueprint for extending transparency and auditability to a broader class of action-oriented systems. Consequently, it may be applicable to environments in which human judgment is vulnerable to fatigue, status quo bias, or desensitization under sustained cognitive load.
  • 1.
    Clinical Triage and High-Dependency Care: In emergency and critical-care settings, clinicians operate under persistent cognitive load, where subtle signs of deterioration may be overlooked due to desensitization or alarm fatigue. A prescriptive auditor could operate over continuous indicators of physiological state evolution—analogous to the performance signals used in this study—to surface latent risk trajectories that remain within nominal ranges but exhibit adverse momentum. Rather than producing diagnostic predictions, the system would prescribe prioritized re-evaluation, supporting timely intervention before irreversible deterioration occurs.
  • 2.
    Financial Risk Management (Trading Desks): In financial decision-making, human operators are known to exhibit systematic biases such as sunk cost effects and loss aversion, leading to prolonged exposure to unfavorable positions. A prescriptive auditing system could continuously evaluate the alignment between risk exposure and market dynamics, issuing normative recommendations when exposure exceeds logic-based thresholds. In this sense, the system functions not as a predictor of market outcomes, but as an auditor of decision states under volatility.
  • 3.
    Football Player and Tactical Management: In professional football, decisions regarding substitutions and tactical adjustments are often influenced by human biases, including status quo bias and outcome bias, which can delay or distort interventions. A prescriptive AI framework could be applied to evaluate decision-relevant states in real time, potentially decoupling recommended actions from match outcomes and mitigating reliance on salient events that may mislead human judgment. By auditing latent risk states and providing interpretable, contestable recommendations, such a system might support interventions—such as substitutions or role adjustments—based on objective assessments of performance deterioration rather than heuristic or retrospective outcome evaluation. The “substitution” logic instantiated in this context may generalize to proactive, bias-resistant management of player capacity and team performance, and could potentially inform decision-making in other high-stakes, real-time human-in-the-loop domains.
Across these domains, the core function of the system remains invariant: to decouple the evaluation of the decision state at time $t$ from stochastic outcomes at time $t+n$, thereby mitigating outcome bias in high-variance, high-stakes environments. To operationalize and rigorously stress-test this prescriptive paradigm, we deliberately select elite professional football as the empirical domain. This choice is methodological rather than domain-driven: elite football constitutes an adversarial natural laboratory in which decisions are time-critical, irreversible, and subject to continuous public scrutiny, while outcome signals are sparse, delayed, and dominated by stochastic variance. These conditions systematically amplify the very cognitive and epistemic failure modes—such as outcome bias, salience-driven miscalibration, and status quo inertia—that prescriptive auditing is designed to expose. Moreover, the high temporal resolution of in-game decision windows enables fine-grained longitudinal analysis of decision states under evolving uncertainty, a property rarely attainable in controlled laboratory settings. Consequently, demonstrating epistemic coherence, auditability, and contestability under the extreme variance and institutional pressures of elite football provides a stringent stress test for the proposed framework and supports its applicability, a fortiori, to other high-stakes human-in-the-loop decision domains.
5 Methodology
5.1 Baseline Limitations and Methodological Motivation
The academic literature on soccer analytics provides the theoretical foundation for this work, situated at the intersection of performance evaluation and tactical decision modeling.
A central methodological challenge in this domain is quantifying player performance. The PlayeRank framework, proposed by Pappalardo et al. [30], defines a multidimensional, role-aware metric that combines individual technical actions with influence on the passing network, validated against professional scout assessments. The authors explicitly aim to reduce biases related to playing time, normalizing performance vectors by the number of actions to “avoid biases due to play time.”
However, although normalization occurs at the level of the action vector, the final match score is computed as a weighted cumulative sum of contributions. When inspected over the temporal progression of a match, this formulation tends to produce strictly increasing trajectories, effectively reintroducing a play-time exposure bias that favors players who remain longer on the field, even if their per-minute contribution rates are similar. This monotonic behavior introduces a methodological limitation for temporal or comparative analyses. Each additional action contributes positively to the cumulative sum—regardless of qualitative impact—so players with identical per-minute performance profiles diverge in total score purely as a function of playing time. While this approach is suitable for match-level summaries, it becomes problematic when the goal is to characterize relative influence or detect shifts in performance dynamics.
The strictly increasing nature of the cumulative score also prevents the representation of performance decay, momentum reversals, or tactical suppression: a player whose influence declines due to fatigue or tactical adjustments will still exhibit an artificially monotonic performance trajectory. Therefore, while PlayeRank’s cumulative formulation is adequate for full-match evaluation, it is insufficient for temporal segmentation. A temporally meaningful metric must reflect evolving game states, substitutions, and local contributions, rather than indiscriminately accumulating actions.
Recent work supports this perspective. Schmidt et al. [42] demonstrate that player influence fluctuates meaningfully over a match due to tactical reconfiguration, opponent pressure, and physical factors, advocating for temporal windowing to capture such dynamics. Following this insight, the present work also adopts 5-minute temporal slices, but advances the methodology by adding a role-aware percentile cumulative mean across slices. This modification preserves interpretability while enabling direct comparability between players with different playing times, eliminating residual exposure bias and allowing the performance curve to reflect true temporal variability.
A second methodological challenge concerns the interpretation of imprecise or uncertain information in decision-making contexts. The literature highlights fuzzy logic as an effective tool for this purpose. For instance, Huarachi-Macuri et al. LACCEI [20] demonstrate that fuzzy systems are suitable for transforming quantitative descriptors into interpretable linguistic assessments (e.g., Stamina, Agility). Within tactical modeling, Marliere [23] uses fuzzy control systems to arbitrate between tactical states under uncertainty. These contributions motivate the role of fuzzy logic in the present system: integrating performance indicators and contextual factors into a Substitution Priority score in a manner compatible with the inherent uncertainty of in-game decision making.
The decision to make substitutions in soccer remains an intrinsically human domain, hardly replicable by machine learning (ML) algorithms. While ML can analyze historical performance data, a coach’s choice involves a complex assessment of contextual, tactical, and psychological factors that are not easily quantifiable—such as team morale, perceived fatigue, reading the opponent’s strategy, and game momentum. Attempts to predict substitutions using ML, while informative, reveal a clear ceiling in their ability to capture this complexity. A recent study by Mohandas et al. [26], for example, analyzed a large dataset of 51,738 substitutions using multiple algorithms, including Random Forest, SVM, and Decision Trees. Even with this wealth of data, the best-performing model (Random Forest) achieved a maximum accuracy of just over 70%. This gap of approximately 30% is significant, suggesting that the final decision is not purely predictive but rather adaptive, relying on subjective human judgment that historical data alone cannot capture. Consequently, models attempting to predict the optimal substitution are conceptually limited. Alternative approaches, such as fuzzy logic-based decision support systems, appear more suitable, as they are designed not to replace but to support the coach’s judgment, managing the inherent uncertainty and qualitative factors of the match.
Finally, the choice of an intrinsically interpretable methodology is not merely a design preference, but a methodological necessity in high-stakes decision-making contexts. As explicitly argued by Rudin [39], post-hoc explanations of black-box models are fundamentally inadequate when decisions have significant consequences, as they obscure the true decision logic and may provide misleading justifications. In domains such as elite soccer, where substitution decisions carry substantial sporting and financial impact, decision support systems must expose their reasoning process in a transparent and auditable manner. This perspective reinforces the limitations of purely predictive machine learning approaches discussed above and provides a principled justification for abandoning black-box optimization in favor of symbolic reasoning. Accordingly, the present work adopts fuzzy logic as the core inference mechanism—rather than as an explanatory add-on—ensuring that substitution recommendations are interpretable, accountable, and directly aligned with domain knowledge and human tactical judgment [39].
This perspective finds foundational support in the framework of logical dynamics proposed by van Benthem [48], which conceptualizes rational agency as a continuous process of information update rather than static optimization. In this context, player substitutions serve as a particularly clear instance of epistemic actions [47]: they do not merely alter the physical parameters of the match, but fundamentally restructure the space of plausible future game states under severe informational constraints. At the moment of decision, the coach operates at the boundary between “hard facts” (e.g., scoreline, remaining time, substitution limits) and “soft information” (e.g., perceived fatigue, tactical momentum). Once executed, a substitution irreversibly fixes certain constraints while reweighting the plausibilities regarding future play. Consequently, decision quality cannot be assessed solely by ex-post outcomes, but by whether the action constituted a rational epistemic update given the information available at time $t$. This framing aligns the proposed system with a dynamic view of rationality—where correctness is understood as the capacity for informed correction under uncertainty—establishing its role not as a predictor of future success, but as an auditor of the epistemic transition induced by the substitution.
Crucially, the proposed approach is hybrid in a functional rather than predictive sense. The statistical components of the system are not employed to forecast future match outcomes or to optimize predictive accuracy. Instead, they serve exclusively to construct a temporally normalized, role-aware representation of the current decision state. The symbolic layer does not refine or correct predictions; it operates as an external normative auditor, evaluating whether continued action is justified given the evidence available at time $t$. In this architecture, hybridization exists to support interpretability, decision auditing, and accountability—not predictive performance.
5.2 Description of the Problem
5.2.1 Problem Formulation
Tactical decision-making in elite soccer operates under high uncertainty and significant financial stakes [14, 12]. Despite the proliferation of granular data, the specific process of substitution decisions remains predominantly intuitive or reliant on descriptive statistics that fail to capture real-time performance decay. This creates two distinct problems: the Metric Exposure Bias and the Predictive Ceiling.
The Metric Exposure Bias: Existing frameworks for player evaluation, such as the original PlayeRank [30], typically utilize cumulative sum formulations. While effective for post-match rankings, these metrics introduce a temporal bias during live games: a player’s score monotonically increases with playing time, regardless of their minute-by-minute efficiency. This mathematical structure masks declining performance, as a fatigued player performing poorly in the $80^{th}$ minute may still have a higher total score than a high-impact substitute, rendering standard metrics insufficient for real-time substitution decisions.
The Predictive Ceiling of Machine Learning: Current computational approaches to substitutions rely heavily on Supervised Machine Learning (SML) models trained on historical data. As noted by [26], these models achieve a prediction accuracy plateau of approximately 70%. This ”ceiling” exists because SML models are designed to mimic human behavior, thereby learning and replicating the cognitive biases of coaches (e.g., status quo bias or delaying defensive changes). Consequently, these models validate historical decisions rather than optimizing future outcomes, failing to identify necessary substitutions that deviate from conservative human norms.
Therefore, the core problem this study addresses is the lack of an objective, prescriptive auditing tool that can quantify intra-match performance decay without exposure bias and signal tactical risks independently of historical human tendencies.
5.2.2 Research Hypotheses
To address the formulated problem, this study tests the following hypotheses:
  • 1.
    H1 (Metric Hypothesis): Reformulating performance evaluation from a cumulative sum to a Cumulative Mean with Role-Aware Normalization effectively eliminates exposure bias, allowing the system to detect negative momentum and performance drops that are mathematically invisible in additive models.
  • 2.
    H2 (Prescriptive Auditing Hypothesis): A prescriptive decision support system, acting as an objective auditor of performance and contextual variables (e.g., performance trends, fatigue, disciplinary exposure), can identify high-risk tactical scenarios—such as latent defensive liabilities—earlier and more reliably than human intuition or data-driven predictive models trained on historical behavior.
  • 3.
    H3 (Contextual Modulation Hypothesis): The modulation of disciplinary risk by tactical role (e.g., weighting a yellow card more heavily for a defender than a forward) significantly alters the substitution priority score, aligning the system’s output with expert tactical consensus in critical defensive scenarios.
  • 4.
    H4 (Prescriptive AI Formalization Hypothesis): The empirical validation of a prescriptive, auditing-based decision support system in a high-stakes environment supports the formalization of a distinct Prescriptive AI paradigm, in which the system objective is not to predict human decisions, but to support normative, risk-aware judgment through explicit reasoning and accountability.
5.3 Dataset
The database selected for this project is the ”Soccer match event dataset”, a detailed public repository of soccer match events [31]. The choice of this dataset is based on its high granularity. The dataset details player-level actions and aggregated performance metrics, allowing for the modeling of in-game situations, which is an essential requirement for our system. We used the complete dataset available on Kaggle [31], which comprises 27 interrelated CSV tables. They cover seven main competitions and can be grouped into five logical categories:
  • 1.
    Match Data (matches_*.csv): Contains information about the games, such as dates, lineups, substitutions, results, and tactical formations.
  • 2.
    Event Data (events_*.csv): The core of the dataset, recording millions of individual actions on the field.
  • 3.
    Entity Data (players.csv, teams.csv, etc.): Dimensional tables with demographic and static data.
  • 4.
    Performance Metrics (playerank.csv): A pre-processed file that provides the playerankScore, a multidimensional and role-aware performance evaluation metric.
  • 5.
    Dictionaries (tags2name.csv, eventid2name.csv): Metadata that translate event and tag IDs into readable descriptions (e.g., Tag 1702 = ’yellow_card’).
The choice of playerankScore as our main ”Performance” input is a central methodological decision. As proposed by Pappalardo et al. [30], the PlayeRank framework was developed to solve the absence of a consolidated and universally accepted metric for evaluating player performance. The playerankScore is a metric derived from millions of game events that, according to the authors, surpasses other metrics when compared with assessments from professional scouts. Therefore, instead of trying to model performance from raw events, we adopted the playerankScore as an already validated and academically robust representation of a player’s performance in a match.
5.4 Data Integration and Pre-processing
The pre-processing pipeline transforms heterogeneous soccer event logs into a unified temporal representation of player performance suitable for fuzzy inference. The process unfolds in three sequential stages: temporal performance computation, contextual and demographic enrichment, and role-aware normalization. Each stage progressively increases the dataset’s semantic density and interpretability.
The objective of this pipeline is to transform the raw multi-source event data (events, matches, players) into a single structured and temporal dataset tailored for the Fuzzy Logic Decision System. Unlike the original PlayeRank framework Pappalardo et al. [30], which aggregates player performance per match, our system requires a fine-grained temporal perspective of the player’s performance evolution within a match. To achieve this, we implemented a three-stage processing pipeline that generates a dense temporal dataset, where each record represents the state of a player in a 5-minute (300-second) interval of play.
5.4.1 Stage 1: Temporal Performance Metric Generation
The first stage (playerankdatasetV4.py) processes the raw event logs (events_*.csv) and converts each action’s timestamp (eventSec, matchPeriod) into an absolute measure of seconds from the start of the match (total_seconds). The match timeline is then discretized into 5-minute intervals, associating each event with its corresponding temporal slice. The following variables are created in this stage:
  • 1.
    matchId, teamId, playerId: extracted directly from the event logs, providing the hierarchical identifiers for each observation.
  • 2.
    cartao_amarelo: a binary state indicator (using expanding().max()) equal to 1 if the player received a yellow card at any point up to the current slice.
Within each interval, two complementary components are calculated:
  • 1.
    Technical Score (score_tecnico_fatia): the aggregation of event-based actions (passes, shots, duels, interceptions) using the same weights of the original PlayeRank framework. Events are mapped to technical dimensions and normalized within the team context.
  • 2.
    Network Score (score_rede_fatia): computed using a directed pass graph per team and interval. Following the original PlayeRank framework, the Eigenvector Centrality of each player is used to quantify their structural influence on ball circulation.
These two components are linearly combined to form the primary performance indicator for each temporal slice:
$\mathrm{playerank\_fatia\_raw}=(1-\alpha_{\mathrm{net}})\cdot\mathrm{score\_tecnico\_fatia}+\alpha_{\mathrm{net}}\cdot\mathrm{score\_rede\_fatia}$
where $\alpha_{\mathrm{net}}=0.2$ controls the relative contribution of the network-based component. This configuration was selected to balance structural influence and technical efficiency, aligning with the sensitivity ranges reported in Pappalardo et al. [30].
Subsequently, to ensure metric interpretability independent of absolute match intensity, the raw score is normalized against the historical distribution of the agent’s specific tactical role. This transformation converts the unbounded scalar playerank_fatia_raw into the Role-Aware Percentile (playerank_fatia_percentil), ensuring that a ”High” score represents the same statistical rarity for a Defender as it does for a Forward.
Cumulative Metric Redefinition.
To capture performance evolution through time, we define the cumulative mean score:
$\mathrm{playerank\_acumulativo\_media\_percentil}\ \ \textbf{=}\ \ \frac{1}{t}\sum_{i=1}^{t}\mathrm{playerank\_fatia\_percentil}^{(i)}$
This cumulative formulation departs from the additive model of the original PlayeRank, ensuring that longer playing times do not artificially inflate the total performance. The metric represents the player’s average contribution up to time $t$, maintaining comparability among players with different play durations.
5.4.2 Stage 2: Contextual and Demographic Enrichment
The second stage (expandirdataset.py) enhances the temporal dataset with contextual, situational, and demographic variables by joining multiple data sources.
  • 1.
    player_age: calculated by merging player metadata (players.csv) with match dates and computing the player’s age on match day.
  • 2.
    player_position: extracted from the player metadata, representing the tactical role (goalkeeper, defender, midfielder, forward).
  • 3.
    goals_scored: identified from events where eventName = "Shot" and subEventName = "Goal" and accumulated over time.
  • 4.
    assists: extracted using Wyscout tag id = 302 and accumulated over time using groupby([’matchId’, ’playerId’]).cumsum().
  • 5.
    momentum_rate: a measure of short-term performance trend, calculated as the difference between the current($t$) and the previous(t1t-1) 5-minute playerank_acumulativo_media_percentil slices.
This enrichment process yields a semantically rich dataset combining technical, contextual, and demographic aspects for each player’s temporal trace.
5.4.3 Stage 3: Final Cleaning and Role-Aware Normalization
The final stage (limpezafinal.py) standardizes the temporal data and recalculates cumulative metrics for stability and accuracy. The following operations are performed:
  • 1.
    removal of out-of-field periods: Periods where it was identified that the player was substituted, or entered as a substitute, were removed to ensure that the metric remained accurate, not applied to goalkeepers.
  • 2.
    Temporal Harmonization: The temporal variables are standardized. Tempo_Partida is set to represent the *end* of each 5-minute slice (e.g., 5, 10, 15…), and minutes_played is recalculated using cumcount() to ensure a precise cumulative sum of on-field time.
  • 3.
    Recalculation of Cumulative Means: The variable
    playerank_acumulativo_media_raw is recalculated using a robust expanding().mean() to ensure that averages reflect the final, harmonized temporal trace.
  • 4.
    Percentile Normalization: two percentile-based variables are computed within each (matchId, position) group. Note: this normalization uses the raw position column (from the original data source) to group players by tactical role, ensuring methodological consistency:
    $\mathrm{playerank\_fatia\_percentil}$ $=\mathrm{rank}_{\mathrm{pct}}(\mathrm{playerank\_fatia\_raw})$
    $\mathrm{playerank\_acumulativo\_media\_percentil}$ $=\mathrm{rank}_{\mathrm{pct}}(\mathrm{playerank\_acumulativo\_media\_raw})$
This role-aware normalization ensures that performance is assessed relative to tactical peers, preserving fairness and interpretability in subsequent fuzzy inference.
Final Dataset Variables.
The resulting dataset (Dataset_limpo_final.csv) contains the following variables:
  • 1.
    Identifiers: matchId, playerId, teamId.
  • 2.
    Temporal Dimensions: Tempo_Partida, minutes_played.
  • 3.
    Performance Metrics: playerank_fatia_raw, playerank_acumulativo_media_raw, score_tecnico_fatia, score_rede_fatia.
  • 4.
    Normalized Metrics: playerank_fatia_percentil,
    playerank_acumulativo_media_percentil.
  • 5.
    Contextual Factors: momentum_rate, cartao_amarelo.
  • 6.
    Demographics: player_age, player_position.
  • 7.
    Offensive Contributions: goals_scored, assists.
  • 8.
    Auxiliary: position, a placeholder variable maintained for compatibility.
Each record corresponds to a 5-minute temporal snapshot of an individual player’s on-field performance, integrating technical, contextual, and demographic information in a consistent temporal framework.
5.4.4 Data Validation
Before integration into the fuzzy system, consistency checks were performed. The recalculation of playerank_acumulativo_media_raw in the final stage (Stage 3) ensures that all cumulative metrics are based on the robust expanding().mean() operator applied to the complete and harmonized temporal trace of each player. This guarantees that all cumulative scores represent valid in-game performance dynamics and that the fuzzy system operates on stable, correctly aggregated data. The experimental design yields high statistical density: the inference engine executes approximately 360 distinct decision audits (2 teams × 10 outfield players × 18 temporal windows). This granular approach allows us to evaluate the system’s stability across diverse game states (winning/losing, high/low fatigue) within a controlled adversarial environment, ensuring internal validity beyond a simple match-outcome correlation
5.5 Exploratory Data Analysis
The Exploratory Data Analysis (EDA) was conducted on the final integrated dataset, consisting of 805,146 temporal observations, 3,035 unique players, and 1,941 matches. The analysis aimed to (i) characterize the statistical properties of the input variables used in the fuzzy inference system, (ii) validate soccer-specific behavioral hypotheses (e.g., positional risk asymmetry), and (iii) provide empirical grounding for the design of the fuzzy universes of discourse.
5.5.1 Descriptive Overview
The dataset preserves a balanced representation of real-world soccer demographics and tactical structures. Table 3 summarizes the player distribution and key descriptive statistics.
Table 3: Dataset Summary and Descriptive Statistics
Category Value Metric Notes
Total observations 805,146 intervals 5-minute resolution
Unique players 3,035 across 7 competitions
Matches analyzed 1,941 complete event coverage
Mean/Median performance 0.501 std = 0.289 centered around 0.5
(playerank...percentil) 0.501 symmetrical distribution
Player positions 4 roles role-aware normalization applied
   Goalkeeper 233 (5.8%)
   Defender 1,032 (25.8%)
   Midfielder 1,120 (28.0%)
   Forward 650 (16.3%)
This composition closely mirrors elite-level soccer distributions, where midfielders and defenders together comprise over half of the active roster, reflecting the denser tactical occupation of central zones.
5.5.2 Univariate Distributions
Player Age.
The age distribution (Figure 3) follows an approximately median centered around 26.0 years, with a range between 15.0 and 45.0 years. The kurtosis (2.91) and near-zero skewness confirm symmetry. This validates the fuzzy partitioning into Young, Peak, and Veteran sets, with smooth transitions around the mean.
Figure 3: Distribution of player ages in the dataset.
Performance (playerank_acumulativo_media_percentil)
The per-slice performance scores exhibit a bounded, quasi-normal distribution centered around 0.5, with a standard deviation of 0.289. These characteristics make the metric well-suited for fuzzy modeling, where zero-centered or bounded scales allow intuitive linguistic partitioning (Low, Medium, High) without rescaling distortions.
5.5.3 Role-Based Event Distributions
To further interpret positional behavior, the dataset was aggregated by player position to estimate empirical probabilities of key in-game events (goals, assists, and yellow cards).
The empirical distributions align with established tactical expectations:
  • 1.
    Goal Rate: Forwards lead with a 19.8% probability of scoring per match, compared to 7.7% for midfielders and 3.4% for defenders. Goalkeepers are near zero, validating the offensive gradient embedded in the player_position variable.
  • 1.
    Assist Rate: Forwards exhibit the highest assist probability (7.5%), reflecting their creative and distributive role in ball progression. Midfielders follow closely (6.8%), while defenders contribute the least(3.6%).
  • 1.
    Disciplinary Risk: Defenders show the highest yellow card rate (17.2%), followed by midfielders (14.8%) and forwards (10.1%), reinforcing the need for contextual fuzzy rules that modulate card risk according to tactical role.
5.5.4 Variable Definition, Independence and Correlations
Variable Definition and Selection
From the final integrated dataset (Dataset_limpo _final.csv), nine variables were selected for exploratory analysis and subsequent use as inputs in the fuzzy decision model. These variables represent the main dimensions of in-game player state: performance, fatigue, disciplinary risk, offensive contribution, and tactical context. Table 4 summarizes the selected features, their measurement scale, and their conceptual role within the system.
Table 4: Variables selected for exploratory analysis and fuzzy modeling.
Dimension Variable (ID) Mathematical Set Description and Relevance Technical Performance P_cum $0,1]\subset\mathbb{R}$ Cumulative performance percentile per match and position; main indicator of global efficiency. Fatigue Min_played $0,100]\subset\mathbb{R}$ Total minutes played at each interval; proxy for physical wear. Disciplinary Risk Card_Y $\{0,1\}\subset\mathbb{Z}$ Indicates if player received a yellow card; higher risk increases substitution priority. Form Trend Momentum $1,1]\subset\mathbb{R}$ Short-term rate of change in performance (3-interval moving average). Negative values denote declining form. Contextual Fatigue (Age) Age $15,45]\subset\mathbb{N}$ Player’s age on match day; modulates fatigue effects. Tactical Role Position $Categorical$ Role label; used for role-aware normalization. Offensive Contribution (Goals) Goals $\mathbb{N}$ Cumulative goals scored; high values indicate offensive importance. Offensive Contribution (Assists) Assists $\mathbb{N}$ Cumulative assists; complements goal contribution.
Labels: P_cum = playerank_acumulativo_media_percentil, Min_played = minutes_played, Card_Y = cartao_amarelo, Momentum = momentum_rate, Age = player_age, Position = player_position, Goals = goals_scored, Assists = assists,
These variables were derived directly from the processed PlayeRank-based dataset, ensuring temporal coherence across 5-minute intervals. The combination of continuous, discrete, and categorical features provides a multidimensional view of player state suitable for fuzzy inference. In the next subsection, we assess the independence and correlation among these variables to confirm their suitability for use as fuzzy inputs.
Spearman Correlation Matrix
Figure 4 reports the Spearman correlation matrix computed over all input variables adopted in the fuzzy inference system. The pairwise correlations are consistently weak, with magnitudes predominantly below $|\rho|<0.20$. Such a pattern indicates that the variables do not exhibit meaningful monotonic dependence.
Figure 4: Spearman correlation matrix among the fuzzy-system input variables. The weak correlations indicate low redundancy and support the assumption of variable independence in fuzzy inference.
This low degree of association is desirable in the context of fuzzy inference. Since the rule base relies on linguistic terms and membership functions that encode distinct semantic dimensions of player performance, strongly correlated variables would introduce redundancy and reduce the discriminative power of the fuzzy rules. Conversely, the weak correlations observed here suggest that each variable contributes independent information to the inference process, thereby supporting a more expressive and interpretable fuzzy model.
5.6 Fuzzy Control System
Based on the theoretical foundation and the need for dynamic real-time evaluation, we designed the Fuzzy Control System (FCS) to function as a Contextual Modifier. Unlike traditional systems that calculate a raw output from zero, this architecture calculates a correction factor applied to a baseline performance metric. The system was implemented using the scikit-fuzzy library in Python, utilizing the Mamdani architecture.
5.6.1 Definition of Fuzzy Variables
The system architecture has been expanded to process high-dimensional match data. It comprises eight input variables (antecedents) and one output variable (consequent). The universes of discourse were calibrated based on the range of values observed in the dataset and domain constraints (e.g., match duration).
Antecedents (Inputs)
  1. 1.
    Cumulative Performance (P_cum): Represents the player’s average percentile performance up to the current moment.
    • (a)
      Universe: [0.0, 1.0].
    • (b)
      Sets: VeryLow, Low, Medium, High, VeryHigh.
  2. 2.
    Momentum (Momentum): Rate of change in performance, indicating if the player is improving or declining.
    • (a)
      Universe: [-1.0, 1.0].
    • (b)
      Sets: Falling, Stable, Rising.
  3. 3.
    Fatigue (Min_played): Minutes played in the match.
    • (a)
      Universe: [0, 100] minutes.
    • (b)
      Sets: Low (0-45’), Medium (40-80’), High (70-100’).
  4. 4.
    Age (Age): Player’s chronological age.
    • (a)
      Universe: [15, 45] years.
    • (b)
      Sets: Young, Peak, Veteran.
  5. 5.
    Match Events (Categorical/Integer):
    • (a)
      Card_Y: Yellow card status [0, 1] (Yes).
    • (b)
      Goals: Goals scored [0, 10] (None, Some, Many).
    • (c)
      Assists: Assists provided [0, 10] (None, Some, Many).
  6. 6.
    Positional Context (Binary):
    • (a)
      Variables: is_Defender, is_Midfielder, is_Forward.
    • (b)
      Sets: Yes (indicating the player’s role to activate specific rules).
Consequent (Output)
  • 1.
    Modifier Value (Modifier_Value): A correction factor ranging from negative (protection/priority reduction) to positive (urgency/priority increase).
    • (a)
      Universe: [-100, 100].
    • (b)
      Granularity: 9 sets ranging from VLN (Very Large Negative, -100) to LP_70 (Large Positive, +70).
    • (c)
      Zero represents no adjustment.
5.6.2 Membership Functions
The membership functions (MFs) mix Trapezoidal (trapmf) and Triangular (trimf) shapes to capture specific tactical thresholds. Key definitions from the implementation include:
  • 1.
    P_cum: Low uses a trapezoid [0, 0, 0.10, 0.35] to capture distinctively poor performance, while High starts at 0.65.
  • 2.
    Momentum: Falling is defined strictly in the negative range [-1.0, -1.0, -0.03, -0.01], ensuring only genuine performance drops trigger the logic.
  • 3.
    Fatigue: Overlapping sets allow smooth transitions. High fatigue begins notably at 70 minutes, aligning with common substitution windows.
  • 4.
    Contextual Events: Variables like is_Defender or Card_Y use pseudo-binary trapezoids (e.g., [0.5, 1, 1, 1.5]) to function as logical switches within the fuzzy inference engine.
5.6.3 Rule Base
The rule base was streamlined to remove noise and focus on Intensifying or attenuating the player’s performance using other contextual events as variables. The system avoids oscillation by ignoring minor fluctuations and focuses on critical states. The logic integrates position and stats directly.
Table 5: Key Fuzzy Rules for Substitution Priority
ID Rule Logic R01 – Untouchable Star IF (P_cum is High/VeryHigh) THEN (Modifier = Negative). R02a – Critical Fatigue IF (P_cum Low/VeryLow) AND (Min High) THEN (Modifier = Large Positive). R02b – Early Fatigue IF (P_cum Low/VeryLow) AND (Min Medium) THEN (Modifier = Medium Positive). R03 – Defensive Risk IF (Defender AND YellowCard Yes) THEN (Modifier = Medium Positive). R04 – Rapid Decline IF (P_cum Low/VeryLow AND Momentum Falling) THEN (Modifier = Large Positive). R07 – Positive Momentum IF (P_cum High/VeryHigh AND Momentum Rising) THEN (Modifier = Medium Negative). R08 – Ineffective Forward IF (Forward AND P_cum Low AND Goals None) THEN (Modifier = Large Positive). R09 – Striker Under Pressure (1) IF (Forward AND P_cum Low/Med AND Momentum Falling) THEN (Modifier = VeryLarge Positive). (2) IF (Forward AND P_cum Low/Med AND Min High) THEN (Modifier = VeryLarge Positive). R10 – Invisible Playmaker IF (Midfielder AND P_cum Low AND Assists None) THEN (Modifier = Med–Large Positive). R11 – Creator Bonus IF (Assists Some OR Goals Some) THEN (Modifier = Medium Negative). IF (Assists Many OR Goals Many) THEN (Modifier = Large Negative). R12 – Veteran Fatigue IF (Age Veteran AND Min High) THEN (Modifier = Medium Positive). R13 – Young Talent Protection IF (Age Young AND (Goals Some/Many OR Assists Some/Many)) THEN (Modifier = VeryLarge Negative). R14 – Goal Protection IF (Goals Some) THEN (Modifier = Medium Negative). IF (Goals Many) THEN (Modifier = Large Negative). R15 – Neutral State IF (P_cum Medium AND Momentum Stable) THEN (Modifier = Zero).
5.6.4 Inference and Final Calculation
The system employs a hybrid calculation model. The Fuzzy Inference System (FIS) computes the Modifier_Value using the Centroid method. However, this value is not the final priority. The final substitution priority $P_{final}$ is calculated by applying the modifier to a baseline inverse of the cumulative performance, scaled by a factor $\alpha$:
$\text{Baseline}=100\times(1.0-P_{\text{cum}})$ (1)
$P_{\text{final}}=\text{clip}\left(\text{Baseline}+(\text{Modifier}\times\alpha),\ 0,\ 100\right)$ (2)
Where $\alpha=0.25$ in the current version. This ensures that the fuzzy logic acts as a ”tuner” that intensifies or attenuates the urgency based on tactical context (e.g., a yellow card or a goal scored), rather than overriding the player’s core performance metric entirely. This methodology allows the coach’s in field knowledge to be faithfully captured by the system, adding significant weight to the player’s performance on the field.
6 Results
To evaluate the utility of prescriptive auditing in high-stakes environments, we contrast the proposed Fuzzy Control System (FCS) against state-of-the-art predictive baselines and subject it to a rigorous stress-test using real-world data from elite soccer.
6.1 Overcoming the Predictive Ceiling: Prescription vs. Prediction
Recent benchmarks in soccer analytics reveal a predictive ceiling for substitution modeling. Mohandas et al. [26] analyzed 51,738 substitutions using supervised learning algorithms such as Random Forest and SVM, achieving a maximum accuracy of approximately $70\%$. This plateau indicates that nearly 30% of tactical decisions are influenced by contextual nuances—such as fatigue thresholds or specific game plans—that remain opaque to black-box models.
Supervised learning approaches in this domain suffer from a fundamental normative limitation: they are trained to mimic human experts. If historical data encode risk aversion, delayed reactions, or other systematic biases, the resulting models replicate them rather than correct them. As summarized in Table 6, our proposed framework shifts the objective from minimizing prediction error to maximizing tactical utility, acting as an auditor rather than an imitator.
Theoretical considerations further reinforce this distinction. As established in Theorem 2, constructing or benchmarking against an additional predictive substitution model does not provide a meaningful comparison. Predictive systems, by design, optimize for behavioral imitation or outcome likelihood. Consequently, they evaluate decisions with respect to variables that are epistemically unavailable at the moment of commitment. As formally shown in Section 3.1.2, such systems are inherently incapable of auditing decision quality without reintroducing outcome bias. Any predictive baseline—regardless of architectural sophistication or empirical performance—would therefore inherit the same normative limitation identified in prior work [e.g., 26]: replication of historical human biases rather than their diagnosis.
From this perspective, the relevant scientific question is not whether a stronger predictor could be engineered, but whether prediction is an appropriate objective for evaluating decisions under uncertainty. Our prescriptive framework answers this question in the negative by evaluating choices according to normative coherence with the epistemic state at time $t$. Requiring predictive parity as a validation standard would constitute a category error, conflating behavioral forecasting with normative decision auditing.
Table 6: Structural Comparison: Supervised ML vs. Prescriptive Fuzzy System
Dimension Supervised ML Baseline (e.g., Mohandas et al. [26]) Proposed Prescriptive System
Objective Minimize Prediction Error (Mimic Human) Maximize Tactical Utility (Audit Human)
Handling Uncertainty Probabilistic (Confidence Scores) Possibilistic (Membership Degrees)
Interpretability Low (Feature Importance/Black Box) High (Linguistic Rules)
Outcome Bias High (Dependent on realized results) Decoupled (State-based evaluation)
6.2 The Mathematical Mechanism: Eliminating Exposure Bias
A fundamental prerequisite for effective auditing is the ability to detect performance deterioration in real-time. Traditional metrics typically utilize cumulative sum formulations, where a player’s score monotonically increases with playing time regardless of minute-by-minute efficiency. Mathematically, these additive models act as low-pass filters, masking declining performance because a fatigued player performing poorly in the $80^{th}$ minute may still possess a higher total score than a high-impact substitute.
To resolve this Metric Exposure Bias, we reformulated the evaluation metric from a cumulative sum to a Cumulative Mean with Role-Aware Normalization ($P_{cum}$). By calculating the average contribution up to time $t$ and normalizing it against the historical distribution of the agent’s specific tactical role, our metric functions as a high-pass audit. This transformation allows the performance curve to reflect true temporal variability, exposing negative momentum and performance drops that are mathematically invisible in additive models. This methodological shift is what enables the system to detect the latent risks detailed in the following case studies.
6.3 Stress-Testing Decision Latency: The Brazil vs. Belgium Case
We applied the framework to the high-stakes elimination match between Brazil and Belgium (2018 FIFA World Cup) to observe decision latency and normative disagreement. The system outputs ($P_{final}$) were generated in 5-minute intervals.
Figure 5: Temporal evolution of Substitution Priority for Brazil. High values indicate critical inefficacy or risk. Note the divergence in the case of Fagner (Red Dashed Line) compared to executed substitutions.
6.3.1 Alignment and Early Detection
The system successfully captured decision-relevant deterioration prior to human intervention (Table 7).
  • 1.
    Willian (45’): Priority $P_{final}=72.0$, ranked 4th. Rule R04 (Rapid Drop) flagged a sharp decline in technical actions in the final 20 minutes of the first half. Immediate substitution by the coach confirms offensive ineffectiveness, supported by media reports [11].
  • 2.
    Gabriel Jesus (58’): Priority $P_{final}=99.1$, 2nd highest. Rule R08 (Ineffective Forward) triggered by low cumulative score and no goals/assists. Post-match data: 12 touches in 58 minutes [38], validating tactical isolation.
  • 3.
    Paulinho (73’): Priority $P_{final}=93.1$. System detected ”Early Fatigue” and low participation. External analysis noted minimal offensive impact [13].
6.3.2 The “Fagner Paradox”: Counterfactual Auditing
The most significant finding is the disagreement regarding right-back Fagner. While the human decision-maker retained the player, the system assigned him Maximum Priority (100.0) from minute 45 to 85. This was not a prediction error, but a risk audit. The system aggregated consistently low technical performance with high defensive exposure against opponent Eden Hazard. The realized outcome—Fagner struggling in duels and receiving a yellow card at the 90th minute—validates the system’s “Critical Risk” diagnosis. A predictive ML model trained on this coach’s history would likely have predicted “No Substitution” (True Positive), thereby reinforcing the status quo bias that led to the vulnerability.
Table 7: Model Audit vs. Human Decision (Brazil vs. Belgium)
Player Analysis Slice($t$) $P_{\text{final}}$ Rank Diagnosis Real Decision
Willian 40’-45’ 72.0 4th Rapid Drop Substituted out
Gabriel Jesus 55’-60’ 99.1 2nd Ineffective Forward Substituted out
Fagner 45’-85’ 100.0 1st Critical Risk Not Substituted
Paulinho 65’-73’ 93.1 3rd Early Fatigue Substituted out
6.3.3 The “Lukaku Paradox”: Latency and Event Masking
Applying the system to the opponent (Belgium, Figure 6) highlights the impact of outcome bias on human decision-making.
Figure 6: Temporal evolution for Belgium. Lukaku reaches maximum priority (100.0) significantly earlier than his substitution.
Romelu Lukaku provided an assist at the 31st minute. While this single salient event seemingly justified his presence to the human manager, the system detected a severe drop in engagement immediately after.
  • 1.
    Algorithmic Detection (35’): The system identified Lukaku as the #1 substitution candidate immediately post-assist, due to critically low volume (14 ball touches in 87 minutes).
  • 2.
    Human Latency (+50 min): The coach waited until the 87th minute to substitute him.
This 50-minute latency demonstrates how prescriptive systems can decouple decision quality from “highlight moments,” ensuring consistent performance evaluation (Table 8).
Table 8: Comparative Analysis of Low-Volume Strikers
Player Context FCS Priority Human Response
G. Jesus (BRA) 12 Touches, 0 G/A High (#2) Substituted (58’)
R. Lukaku (BEL) 14 Touches, 1 Assist Max (#1) Delayed Sub (87’)
6.4 Systemic Validation: Longitudinal Patterns
Extending the analysis beyond single instances, we processed a representative subset of 25 World Cup matches—including the complete Brazilian campaign. This broader scope, comprising N = 9,152 decision windows, reveals structural regularities in how the system audits performance.
Technical Anchors vs. Chronic Inefficacy.
The system consistently distinguished between effective stability and passive stagnation. Key players like Coutinho and Marcelo functioned as “Technical Anchors,” maintaining priority scores near zero regardless of match time. Conversely, the system identified “Chronic Inefficacy” in players like Gabriel Jesus across multiple matches, where priority scores escalated without salient errors, supporting the hypothesis that the system detects latent risks invisible to event-based analysis.
Temporal Gap and Anticipation.
A systematic temporal gap was observed between the system’s “Critical” signal ($P_{final}>90$) and human intervention. In cases like Paulinho (vs. Serbia), the system flagged inefficacy several minutes prior to the substitution, quantifying the decision latency inherent in human processing under pressure.
Post-Entry Validation.
By tracking priority scores after a substitution, the system enables counterfactual assessment. Substitutes like Renato Augusto exhibited immediate monotonic decreases in priority (indicating high impact), whereas others (e.g., Fernandinho vs. Switzerland) maintained high risk scores, objectively characterizing the substitution as ineffective.
6.5 Boundary Conditions: The Limits of Performance Auditing
The system’s epistemic boundaries were rigorously tested by the case of Nacer Chadli (Belgium vs. Brazil), who was substituted at the 83rd minute due to an acute injury. This was the only substitution in the analyzed set not anticipated by the FCS ($P_{final}<50.0$).
Far from a systemic failure, this “false negative” provides crucial validation of the framework’s internal logic. Since the injury was a sudden, exogenous event unrelated to prior performance decay or observable fatigue, the system correctly refrained from flagging the player. This confirms that the high priority scores assigned in other cases (e.g., Lukaku, Jesus) were driven by genuine performance degradation, not by a generalized bias toward late-game substitutions. The framework effectively distinguishes between tactical necessity (predictable via data) and force majeure (epistemically inaccessible).
7 Discussion
This study leverages elite sports as a naturalistic laboratory to audit human judgment under uncertainty, revealing systematic divergences between normative risk assessment and expert human behavior. Our primary contribution to human decision science is the quantification of decision latency—the temporal gap between the statistical accumulation of risk and the human reaction to it. By deploying a system designed to decouple decision quality from stochastic outcomes, we provide empirical evidence that human experts in high-stakes environments are subject to structural cognitive biases—specifically status quo bias and outcome bias—which systematically delay necessary interventions.
To rigorize this auditing approach, we formalize the concept of Prescriptive AI:
Definition. Prescriptive AI is a class of intelligent systems whose primary function is to audit, justify, and support human decisions under uncertainty, rather than to automate actions or predict future outcomes.
Unlike standard predictive analytics, this paradigm is defined by five normative objectives:
  • 1.
    A primary function of decision auditing rather than decision automation;
  • 2.
    Explicit reasoning under uncertainty and asymmetric risk;
  • 3.
    Intrinsic interpretability as a structural requirement, not a post-hoc feature [39];
  • 4.
    Contestable recommendations that preserve human agency;
  • 5.
    Evaluation criteria that decouple state assessment from outcome realization, explicitly avoiding outcome bias [4].
These properties impose normative constraints that exclude many systems commonly labeled as prescriptive. While the framework itself is model-agnostic, the system presented here serves as a concrete instantiation of these principles. By applying this approach, we show that the limitation of human experts is not a failure of strategic intent, but a failure of real-time information processing: humans struggle to detect non-salient performance decay (the “boiling frog” effect) until a catastrophic or highly salient event occurs.
7.1 Behavioral Audit: Quantifying Cognitive Biases
The empirical disparities between the prescriptive system and human agents are not random errors, but operationalizations of well-documented cognitive heuristics:
The “Fagner Paradox” as Status Quo Bias.
The disagreement regarding the defensive player Fagner illustrates the status quo bias (or omission bias) under pressure. While the system identified a “Critical Risk” state ($P_{final}=100.0$) based on sustained defensive exposure, the human manager chose inaction. This behavior reflects an asymmetric preference for maintaining the current state, requiring substantially stronger evidence to justify change than to justify persistence. Given the informational state available at time $t$, this inertia is normatively difficult to justify.
The “Lukaku Paradox” as Salience Masking and Outcome Bias.
The case of the opposing striker demonstrates how outcome bias distorts real-time evaluation. A single salient positive event (an assist) masked a severe and prolonged deterioration in engagement (over $50$ minutes of inactivity). We term this phenomenon Salience Masking: the overweighting of vivid, recent outcomes at the expense of cumulative base-rate information. The prescriptive system—immune to the “halo effect” of the assist—identified the decision to maintain the player as a high-risk error well before the human agent reacted.
7.2 Normative Foundations: Decision-Making as Epistemic State Transition
The validity of this auditing approach rests on a dynamic conception of rationality, grounded in the logical dynamics framework of van Benthem [48]. Under this view, rationality is a property of the epistemic update at the moment of commitment ($t$), rather than of the stochastic realization of outcomes at a later time ($t+n$). If decision quality is defined prior to outcome realization, then systems designed to support human judgment must operate strictly on the informational state available at the moment of action. This perspective justifies our rejection of outcome-based evaluation metrics: a decision is not “correct” because the team won, but because it constituted a normatively coherent response to the risk state at time $t$.
Naturally, any prescriptive audit is contingent on the chosen normative model of risk. Alternative normative assumptions would yield different—but still auditable—recommendations, without undermining the core principle of decoupling decision quality from outcome realization.
7.3 Implications for Human Decision-Making
These findings extend classical accounts of Bounded Rationality in dynamic environments. We argue that decision latency is not merely a domain-specific artifact, but a domain-general construct reflecting the cognitive cost required to override heuristic inertia in the presence of uncertainty and time pressure.
  • 1.
    Filtering and Blind Spots: Real-time tactical management requires processing high-dimensional, noisy data streams. To preserve strategic coherence, human agents apply low-pass cognitive filters that attenuate short-term fluctuations. Our results indicate that this adaptive filtering introduces a structural blind spot for gradual but systematic performance decay.
  • 2.
    AI as Cognitive Orthotics: Accordingly, the role of Prescriptive AI shifts from an “oracle” that predicts future states to a cognitive orthotic—a support structure designed to compensate for specific, predictable human limitations. By surfacing latent risk through interpretable signals (e.g., Rule R04: Rapid Drop), the system confronts decision-makers with evidence their heuristic reasoning would otherwise suppress.
7.4 Normative, Epistemic, and Institutional Implications of Prescriptive AI
7.4.1 Normative Scope and Limitations
A prescriptive audit is only as meaningful as the normative assumptions it encodes. The framework proposed here deliberately commits to an explicit model of risk, temporal aggregation, and asymmetric costs, enabling transparent evaluation of decision quality at the moment of commitment. This explicitness, however, also implies that prescriptive recommendations are inherently contingent rather than universally binding. Alternative normative models—reflecting different institutional priorities, ethical trade-offs, or risk tolerances—would yield different but equally auditable prescriptions.
Crucially, this contingency does not weaken the prescriptive paradigm; rather, it constitutes its primary epistemic advantage. By making normative assumptions explicit and contestable, Prescriptive AI avoids the false objectivity often implied by outcome-optimized or imitation-based systems. Disagreement over normative choices is therefore not a failure of the system, but an expected and desirable feature of a transparent decision audit.
7.4.2 Prescriptive Auditing versus Predictive Alignment
It is important to distinguish prescriptive auditing from predictive alignment objectives commonly pursued in contemporary AI systems. Alignment-based approaches seek to minimize divergence between machine outputs and observed human behavior, implicitly treating historical expert decisions as normative ground truth. In contrast, prescriptive auditing explicitly allows—and indeed foregrounds—systematic disagreement between human judgment and normative risk assessment.
The objective of Prescriptive AI is not convergence between human and machine, but diagnosability of decision-making processes. Persistent divergence identifies regions in which human heuristics, institutional pressures, or cognitive biases override normatively justified responses. From this perspective, disagreement is not an error signal to be minimized, but an epistemic signal to be analyzed.
7.4.3 Robustness under Distributional Shift
A central limitation of predictive decision-support systems is their sensitivity to distributional shift. Models trained to forecast outcomes or imitate historical behavior implicitly assume stability in both the environment and expert competence. When these assumptions fail—as they routinely do in high-variance, adversarial domains—predictive performance degrades and interpretability diminishes.
Prescriptive auditing offers a fundamentally different robustness profile. Because recommendations are derived from explicit normative criteria applied to the current epistemic state, their validity does not depend on the stationarity of empirical distributions. While statistical estimates may fluctuate, the criteria governing decision evaluation remain stable. This property is particularly important in dynamic institutional settings, where tactics, incentives, and environmental conditions evolve faster than predictive models can be reliably retrained.
7.4.4 Human–AI Disagreement as an Epistemic Signal
The systematic disagreements observed between the prescriptive system and human decision-makers constitute a primary object of scientific interest. Rather than treating these divergences as isolated errors, the prescriptive framework enables their localization, quantification, and temporal analysis. Disagreement patterns reveal where human agents systematically discount cumulative risk, overweight salient events, or defer action due to heuristic inertia.
This reframing elevates human–AI interaction from an optimization problem to an epistemic diagnostic tool. The goal is not to replace human judgment, but to expose its structural limitations under uncertainty. In this sense, prescriptive systems function less as decision-makers and more as instruments for auditing the rational coherence of human action.
7.4.5 Prescriptive Auditing versus Causal Attribution
Prescriptive auditing must be clearly distinguished from causal inference and counterfactual explanation. The framework does not seek to explain why outcomes occurred, nor to assign causal responsibility to individual decisions. Instead, it evaluates whether a decision was normatively justified given the information available at the moment of commitment.
By operating strictly on epistemic states rather than outcome-generating mechanisms, prescriptive systems avoid retrospective rationalization. A decision is not deemed incorrect because it led to an unfavorable outcome, nor correct because it coincided with success. This separation is essential for maintaining normative coherence in environments dominated by stochastic variance.
7.4.6 Institutional Scalability and Governance
While the present analysis focuses on individual expert decisions, the prescriptive paradigm naturally extends to institutional contexts. Aggregated audit trails enable organizations to identify recurrent blind spots, structural incentives for inaction, and systematic deviations from stated risk policies. Unlike outcome-based evaluations, which conflate skill with luck, prescriptive records permit longitudinal assessment of decision quality independent of stochastic realization.
This property has direct implications for governance in domains subject to public accountability. In settings where decisions must be justified to regulators, stakeholders, or courts, prescriptive auditing provides a defensible record of rational deliberation grounded in the information available at the time of action.
7.5 Ethical Accountability and Epistemic Trust
Finally, we argue that in high-stakes human–AI teaming, interpretability is a pre-condition for legitimacy. Auditing human judgment requires more than predictive accuracy; it requires explanations that enable decision-makers to validate machine assessments against their own reasoning.
Resolving Epistemic Opacity.
In domains subject to intense public scrutiny—such as elite sports, clinical triage, or emergency operations—opaque “black-box” models pose a substantial epistemic risk. When an algorithm recommends a counter-intuitive action (e.g., substituting a star player), a lack of transparency forces stakeholders to choose between blind faith and outright rejection. By providing a transparent audit trail in which each recommendation is traceable to explicit linguistic rules, the proposed fuzzy framework reduces epistemic asymmetry and transforms the interaction from blind obedience into informed deliberation.
The Audit Trail as Institutional Defense.
Beyond decision support, prescriptive systems serve a protective institutional function. In environments vulnerable to suspicion—whether due to financial incentives, reputational risk, or public accountability—a documented audit trail allows decision-makers to demonstrate that their actions (or inaction) were consistent with a normative risk assessment given the information available at time $t$. In this sense, the system functions not merely as a tactical aid, but as an objective witness to the rationality of the decision process.
Ultimately, this work advances prescriptiveness as an ethical design principle. In contexts where decisions are public, consequential, and irreversible, only transparent, auditable, and human-centered systems can achieve the level of trust required to augment—rather than replace—human agency.
7.6 The Prescriptive Paradigm and the Future of AI Alignment
Current debates on AI alignment largely focus on aligning model outputs with human preferences or values, typically through techniques such as reinforcement learning from human feedback (RLHF) or constitutional AI [1, 3]. These approaches implicitly assume that historical human judgments constitute valid ground truth for training objectives.
The Prescriptive AI paradigm challenges this assumption fundamentally. Rather than treating human decisions as labels to be replicated, prescriptive systems treat them as objects of audit. This inversion has profound implications for AI safety and governance:
Alignment as Auditing, Not Imitation.
In domains where human experts exhibit systematic biases—as demonstrated empirically in this work—alignment through imitation perpetuates rather than corrects decision errors. The Imitation Incompleteness Theorem formalizes this limitation: without external normative signals, supervised learning cannot escape the structural biases embedded in training data.
Prescriptive AI reframes alignment as a bidirectional process: rather than aligning AI to human behavior, the system surfaces divergences that allow humans to align their decisions to normative criteria. This constitutes a form of normative bootstrapping, in which human and machine co-evolve toward improved decision quality.
Implications for High-Stakes AI Deployment.
As AI systems are deployed in safety-critical contexts—autonomous vehicles, medical diagnosis, financial regulation—the tension between imitation and prescription becomes acute. A self-driving car trained solely on human driving behavior will reproduce human errors (e.g., delayed reaction to hazards). A prescriptive system, by contrast, evaluates driving decisions against explicit safety criteria, independent of typical human response times.
This distinction suggests that AI alignment objectives must be domain-dependent: in creative or preference-driven tasks, alignment through imitation may be appropriate; in high-stakes, normatively constrained domains, prescriptive auditing is essential.
The Prescriptive-Predictive Frontier.
Future research must delineate the boundary conditions under which each paradigm applies. We conjecture that prescriptive systems are necessary when:
  1. 1.
    Decisions are irreversible with asymmetric consequences,
  2. 2.
    Historical decisions encode systematic biases,
  3. 3.
    Normative criteria can be formalized independently of outcomes,
  4. 4.
    Human accountability requires explicit justification.
Conversely, predictive systems remain appropriate when human preferences constitute valid optimization targets and environmental dynamics are sufficiently stable.
Establishing this frontier formally—potentially through a generalization of Imitation Incompleteness Theorem to multi-agent or multi- objective settings—constitutes a critical research agenda for trustworthy AI.
7.7 Prescriptive AI and Human Expertise: Augmentation, Not Replacement
A recurring concern in human-AI collaboration research is the risk of deskilling: as systems assume cognitive tasks, human experts lose proficiency, creating long-term dependency [7, 2]. Prescriptive AI presents a distinct profile in this regard.
Preserving Deliberative Capacity.
Unlike fully automated systems, prescriptive auditing retains human agency at the point of commitment (Axiom 4: Contestability). The coach can—and frequently does—override algorithmic recommendations. This preserves the exercise of judgment even when augmented by computational support.
Empirical evidence from our case studies supports this: in 3 of 4 substitutions (Willian, Jesus, Paulinho), the human decision aligned with the system. In 1 case (Fagner), the human overrode the system—demonstrating that decision authority remained with the expert. This is fundamentally different from automation, where human input is reduced to supervisory monitoring.
Skill Transformation, Not Atrophy.
Rather than replacing expertise, prescriptive systems transform the skill profile required:
  • 1.
    From pattern recognition to norm evaluation: Experts shift from detecting performance decay (now automated) to evaluating whether system recommendations align with strategic context.
  • 2.
    From reactive to proactive: Alerts surface latent risks before they become salient, enabling anticipatory rather than reactive intervention.
  • 3.
    From intuition to justification: Decisions must be explained (to stakeholders, media, institutions), elevating the role of explicit reasoning over tacit intuition.
This skill transformation is analogous to how calculators shifted mathematical expertise from computation to problem formulation. The cognitive labor remains, but its locus changes.
Risk of Over-Reliance: Automation Bias Redux.
Despite preserving agency, prescriptive systems remain vulnerable to automation bias [32]: humans may defer excessively to algorithmic recommendations, even when contextual factors warrant override.
Our framework mitigates this through:
  1. 1.
    Transparency: Rules and thresholds are inspectable,
  2. 2.
    Contestability: Overrides are explicitly documented,
  3. 3.
    Audit trails: Post-hoc review of decisions vs. recommendations.
Longitudinal field studies are needed to assess whether these safeguards suffice in practice. We hypothesize that over-reliance risk correlates with:
  • 1.
    System accuracy (higher accuracy $\rightarrow$ stronger deference),
  • 2.
    Organizational culture (hierarchical $\rightarrow$ less contestation),
  • 3.
    Temporal pressure (urgency $\rightarrow$ default to algorithm).
The Role of Training.
Effective deployment of prescriptive AI requires training decision-makers in system critique, not system operation. Experts must learn to:
  • 1.
    Identify edge cases where rules may misfire,
  • 2.
    Recognize when contextual factors (invisible to the system) override statistical signals,
  • 3.
    Calibrate confidence in recommendations based on situational uncertainty.
This represents a new pedagogical challenge: teaching humans to collaborate with AI as critical auditors, not passive consumers. Educational programs in medicine, finance, and other high-stakes domains must integrate training in prescriptive system literacy.
8 Acknowledgments
In the spirit of Open Science, the complete source code—including all pre-processing pipelines, fuzzy inference engines, and validation scripts—is publicly available at https://github.com/Pedro-Passos77/AI-Assisted-Substitution-Decisions-A-Fuzzy-Logic-Approach-to-Real-Time-Game-Management . This manuscript constitutes a substantially expanded and revised version of prior research accepted by the Wharton Sports Analytics Journal and by the AAAI 2026 Bridge on LM Reasoning [34, 35]. The present work introduces significant extensions in methodological robustness, theoretical grounding, the proposed decision-making paradigm, and an expanded case study.
We gratefully acknowledge Luca Pappalardo and collaborators for making the Soccer Match Event Dataset publicly available [30], which was essential to this study. The authors declare no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper, neither this research has received any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Generative AI tools (Gemini 3 Pro and ChatGPT 5.2) were used to assist with programming tasks and the structural organization of the manuscript under strict human supervision. The authors retain full responsibility for the final content, interpretations, and normative justification of the proposed framework.
8.1 Limitations and Future Work
While the proposed Fuzzy Control System (FCS) demonstrates substantial utility as a prescriptive auditing tool, it is essential to delineate its epistemic scope and operational boundaries. The system is not designed to perform causal intervention, predict exogenous events, or replace human judgment. Its function is strictly limited to auditing the observable decision space given the information available at time $t$, and should therefore be interpreted as a normative support mechanism rather than an autonomous decision-maker.
From a data perspective, the reliance on minutes_played as a proxy for physical fatigue assumes a quasi-linear accumulation of workload, serving as an estimation rather than a direct physiological measurement. This abstraction does not capture inter-individual physiological differences (e.g., genetics), variations in match intensity such as the volume of high-speed running, or metabolic expenditure under sustained defensive pressure. As a result, the model may underestimate fatigue in high-exposure contexts or overestimate wear in structurally low-intensity tactical systems.
At the modeling level, the fuzzy rule base is statically defined using expert heuristics. While this design choice is central to intrinsic interpretability and contestability, it may limit generalization across divergent tactical philosophies (e.g., high-pressing versus possession-oriented systems) without manual recalibration of membership functions or rule weights ($\alpha$). This reflects a broader trade-off between explainability and adaptive flexibility that is intrinsic to symbolic reasoning systems.
Importantly, the framework does not account for abrupt exogenous events such as acute injuries, referee decisions, or random disruptions, which occur independently of observable fatigue or performance decay. Such events lie outside the epistemic reach of any decision-auditing system grounded in pre-intervention signals.
Future iterations of this framework aim to address these limitations through three complementary research directions:
  1. 1.
    Integration of Biometric Telemetry: Incorporating GPS, accelerometer, and physiological signals to replace temporal proxies with metrics grounded in both internal and external workload, thereby refining fatigue estimation under heterogeneous match conditions.
  2. 2.
    Neuro-Fuzzy Adaptation (ANFIS): Developing an Adaptive Neuro-Fuzzy Inference System that enables dynamic calibration of membership functions and rule weights through supervised learning on historical expert decisions, preserving interpretability while improving contextual adaptability.
  3. 3.
    Tactical Profile Matching: Extending the framework beyond substitution priority to recommend replacement profiles based on detected tactical deficiencies, enabling decision support not only on whether to intervene, but also on how to intervene given available resources.